### v2

# Pragmatic Reasoning improves LLM Code Generation

Abstract
Large Language Models (LLMs) have demonstrated impressive potential in translating natural language (NL) instructions into program code. However, user instructions often contain inherent ambiguities, making it challenging for LLMs to generate code that accurately reflects the user’s true intent. To address this challenge, researchers have proposed to produce multiple candidates of the program code and then rerank them to identify the best solution. In this paper, we propose CodeRSA, a novel code candidate reranking mechanism built upon the Rational Speech Act (RSA) framework, designed to guide LLMs toward more comprehensive pragmatic reasoning about user intent. We evaluate CodeRSA using one of the latest LLMs on a popular code generation dataset. Our experiment results show that CodeRSA consistently outperforms common baselines, surpasses the state-of-the-art approach in most cases, and demonstrates robust overall performance. These findings underscore the effectiveness of integrating pragmatic reasoning into code candidate reranking, offering a promising direction for enhancing code generation quality in LLMs.
Pragmatic Reasoning improves LLM Code Generation
1 Introduction
Recent advances in generative large language models (LLMs) have demonstrated their impressive ability to generate program code from user-provided natural language instructions (Liu et al., 2024b; Coignion et al., 2024). However, given the intrinsic complexities of coding and the potential ambiguities in user input, producing code in a single attempt may fail to explore the vast solution space, overlooking correct or higher-quality solutions (Liu et al., 2024a). A standard practice to address this shortcoming is to sample multiple solutions, which we refer to as “code candidates” (Chen et al., 2021; Brown et al., 2024), and to rerank them. Researchers have proposed various reranking strategies for code candidates, broadly divided into execution-driven and content-driven approaches. Due to the safety-risks associated with execution-driven approaches (Yetiştiren et al., 2023), we here focus on content-driven methods. These evaluate the generated text, often relying on token-level probabilities. For example, Coder reranking scores each candidate based on the cumulative probability of its tokens, sometimes however favoring “degenerate solutions” (generic or repetitive code) with disproportionately high token probabilities (Zhang et al., 2023).
Figure 1: A comparison of our approach CodeRSA (top) compared to the CoderReviewer method (bottom).
When viewing code generation as a communicative process in which an LLM listens to the user’s intentions (Ouyang et al., 2022), Coder reranking evaluates candidate solutions solely from the listener’s perspective. Yet, research on human communication suggests that effective listeners reason about the speaker (who in turn reasons about the listener) Grice (1975). Frank and Goodman (2012; 2016) provided a principled method for quantifying this process based on a probabilistic framework based on game-theoretic notions, called the Rational Speech Act (RSA) framework. Pu et al. (2020, 2024) demonstrated the RSA framework’s effectiveness on program generation for a simple domain, while Schuster et al. (2024) reported negative results on a spreadsheet domain. One aspect that has held back RSA models from scaling up to realistic use cases is the computational overhead (Pu et al., 2024): it requires reasoning about the set of alternative instructions that the speaker could have given and about the set of alternative pieces of code that could solve the problem, which is very computationally expensive. Zhang et al. (2023) therefore proposed CoderReviewer reranking as a simplified scalable approach that simplifies these probability estimation processes over alternatives. However, it comes at the cost of not fully modelling the dialogic, interactive reasoning that can emerge when speaker and listener exchange information.
This work proposes CodeRSA, enabling LLMs to reason as pragmatic listeners and rank code candidates based on the user’s underlying intentions. It addresses the probability estimates for the set of alternative code candidates and alternative utterances via a sampling approach. CodeRSA generates multiple code candidates, and then generates additional instructions for each candidate, forming a set of potential instructions (including the original one), as illustrated in Fig. 1: the literal listener $L_{0}$ estimates the probability of each code candidate given each instruction from the potential instruction set. The pragmatic speaker $S_{1}$ then applies a special normalization on these probabilities to quantify how specifically an instruction fits the generated code. By comparing these pragmatic speaker scores for the original instruction across all candidates, the pragmatic listener identifies the code candidate that most closely matches the original instruction from the speaker’s perspective, finalizing the reranking process (see Fig. 1).
We conducted experiments using CodeRSA with Llama-3-8B-Instruct, one of the latest language models from the Llama family (Grattafiori et al., 2024) on OpenAI’s HumanEval benchmark (Chen et al., 2021). Our experiment results reveal that CodeRSA reliably outperforms the Coder and CoderReviewer reranking methods. Our qualitative analysis illustrates how the CodeRSA reranking enables better candidate selection, promoting a more comprehensive understanding of user intent.
2 Related Work
Natural Language to Code. Previous research has extensively explored generating code from natural language using neural network models (Ling et al., 2016; Rabinovich et al., 2017; Hayati et al., 2018). Recently, large language models (LLMs) have propelled significant advances in this area, driven by the transformer Vaswani (2017) architecture and large-scale pretraining. Their performance on code generation tasks often surpasses that of traditional models, and in many cases even rivals human programmers (Ni et al., 2024; Becker et al., 2023). Moreover, a recent study shows that LLMs also exhibit strong performance in code summarization, effectively translating code snippets into text (Akib et al., 2024).
Diversity Sampling in Code Generation. In a prior study, Chen et al. (2021) found that allowing the model to generate more candidates significantly raises the probability of yielding a correct answer. This practice, which encourages the model to produce a broader range of potential outputs, is often referred to as diversity sampling.
Code Reranking Methods. When LLMs generate multiple code candidates in response to the user instruction, execution-driven reranking methods such as CodeT (Chen et al., 2022) and AgentCoder (Huang et al., 2024) involve running test cases to gauge each candidate’s correctness. Although often effective, execution-driven approach can introduce additional safety risks and practical obstacles, such as the potential for malicious code execution, environment setup overhead, and resource constraints (Yetiştiren et al., 2023; Khoury et al., 2023). In contrast, content-driven reranking methods are far more versatile because they do not rely on execution and are not even confined to coding tasks.
Coder reranking. In prior work, Chen et al. (2021) reranked code candidates by estimating $\text{P}(c\mid i)$, where $c$ denotes the generated code candidate and $i$ denotes the given instruction. This process can also be called Coder reranking because the LLM is a mere Coder that estimates the candidate probability based on the corresponding instruction.
When using an LLM to estimate conditional probabilities, we compute the probability of each token iteratively. For example, in Coder reranking, the model processes candidate’s tokens from left to right: at each step, it calculates the probability of the current token given the instruction and the previously generated tokens, then appends that token to the context before moving on. The product of these sequential probabilities across all tokens yields the overall probability of the code candidate under the given instruction:
$\text{P}(c\mid i_{0})\;=\;\prod_{t=1}^{|c|}\text{P}_{\text{LLM}}\bigl{(}c^{(t)}\mid i_{0},c^{(<t)}\bigr{)},$
where $c^{(t)}$ denotes the token at position $t$ in the sequence $c$, and $c^{(<t)}$ represents the sequence of all tokens before position $t$.
CoderReviewer reranking.
Zhang et al. (2023) added the concept of a reviewer to Coder reranking (i.e., reassessing whether the instruction matches the generated code candidate). Formally, CoderReviewer metric is represented as follows:
CoderReviewer $=\text{P}(c\mid i)\cdot\text{P}(i\mid c)$
    (Coder)      (Reviewer)
The prompt positions of the instruction and code candidates are reversed in the reviewer’s implementation. Thus, the code-generation task is reformulated as an instruction-generation task to calculate the probability of a given instruction. The CoderReviewer metric is also considered a specialized form of maximum mutual information (Li and Jurafsky, 2016), measuring the bidirectional alignment between the generated code and the input instruction.
3 CodeRSA
This section introduces CodeRSA, an approach that employs the Rational Speech Act (RSA) framework to enhance the reranking of candidate code snippets. CodeRSA extends the models proposed by Cohn-Gordon et al. (2019) and Schuster et al. (2024). The core innovation in CodeRSA arises from the pragmatic listener, which is responsible for selecting and reranking code candidates. It does so by imagining how a “pragmatic speaker” would choose an instruction that best distinguishes the intended code among various potential instructions.
A pragmatic speaker evaluates how effectively a potential instruction communicates the intended code candidate. Within the RSA framework, this evaluation process necessitates an explicit calculation of probabilities. However, given the infinite combinations of possible instructions and code candidates, developing a robust and comprehensive probability estimation strategy becomes essential. To address this challenge, CodeRSA generates additional instructions for each candidate, thereby constructing an expanded set of potential instructions (including the original). Finally, by considering the potential instruction set, a pragmatic speaker can select the most appropriate instruction for each code candidate based on the motivation to effectively convey the message to a literal listener.
At the foundation of CodeRSA is the literal listener, who estimates the probability of each code candidate by interpreting the user’s instruction word for word, without inferring deeper speaker intent.
Literal Listener. A literal listener (denoted $L_{0}$) represents the simplest level of reasoning in the RSA framework. It interprets utterances solely according to their literal meaning, without any higher-level pragmatic reasoning. Theoretically, let $c$ be a candidate code and $i$ be an instruction, then:
$\text{P}_{L_{0}}(c\mid i)\propto\llbracket i\rrbracket(c),$
where $\text{P}_{L_{0}}(c\mid i)$ denotes the literal listener’s estimation of probability of candidate $c$ given instruction $i$. The interpretation function $\llbracket i\rrbracket$ maps an instruction $i$ to a function that takes a code candidate $c$ and returns 1 if $c$ correctly implements $i$, and 0 otherwise.
Likewise, most LLMs generate code by conditioning only on the user’s input instruction, which motivates CodeRSA’s use of an LLM to instantiate this literal listener:
$\text{P}_{L_{0}}(c\mid i)=\text{P}_{\text{LLM}}(c\mid i),$
The literal listener provides CodeRSA’s baseline interpretation of user instructions, serving as the foundation for all subsequent reasoning. Note that the Coder reranking can also be considered a literal listener. For convenience, we refer to the logarithm of $\text{P}_{L_{0}}(c\mid i)$ as the Coder score.
Pragmatic Speaker. In the RSA framework, the pragmatic speaker (denoted $S_{1}$) is primarily responsible for determining whether an instruction $i$ effectively conveys the intended meaning of a candidate code $c$ to the literal listener. Formally, a pragmatic speaker can be defined as:
$\text{P}_{S_{1}}(i\mid c)=\frac{\exp\left(\log\text{P}_{L_{0}}(c\mid i)-C(i)\right)}{\sum_{i^{\prime}}\exp\left(\log\text{P}_{L_{0}}\left(c\mid i^{\prime}\right)-C\left(i^{\prime}\right)\right)},$
Here, $\operatorname{C}(i)$ denotes a cost function that quantifies the “expense” of employing a particular instruction. This formula quantifies the pragmatic speaker’s estimation of the probability that a particular instruction will be used to describe a code candidate, as interpreted from the literal listener’s perspective. Note that an idealized RSA approach would consider every possible instruction $i^{\prime}$, which is infeasible in practice. Instead, CodeRSA leverages the fact that ambiguities in the original instruction often surface in the code candidates generated by LLMs, thereby enabling a sample-based estimation.
Given a user-provided instruction $i_{0}$, we request multiple candidate solutions $\{c_{1},\dots,c_{n}\}$ from the LLM. Each candidate $c_{j}$ may capture a slightly different interpretation of $i_{0}$. We then derive one or more instructions from each candidate $c_{j}$, forming a set of potential instructions $\{i_{0},i_{1},\ldots,i_{m}\}$, which we denote by $I$. This approach ensures that we sample relevant alternative instructions directly tied to how the model interprets the original instruction. Potential instructions then constitute a finite sample set that approximates the otherwise infinite instruction space for pragmatic speaker estimation.
To simplify the model and focus on core pragmatic reasoning, we assume a uniform cost for all instructions, which effectively cancels out during normalization. However, a more detailed modeling of the cost function may provide additional insights, a point we further discuss in the Section 6. A pragmatic speaker then can be defined in a simplified form as:
$\text{P}_{S_{1}}(i\mid c)=\frac{\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})},$
In practice, the distribution of $\text{P}_{L_{0}}(c\mid i)$ is often highly skewed, with a few high probability candidates dominating and a large spread in probability estimates. To address this, we here propose to instead work with log probabilities, which in practice brings low probability alternatives closer together and gives more weight to the relative trends among instruction–code pairs. The pragmatic speaker’s preference score for a given instruction is hence quantified as:
$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})}\;,$
$\text{R}_{S_{1}}(i\mid c)\;\in\;(-1,\,0).$
In log space, after normalizing with respect to log probabilities, smallest numbers will now correspond to the events with the highest probability. Therefore, we invert the sign to negative in order to still choose maximal score instructions. Notably, if $c$ is overly generic or appears plausible under multiple instructions, then the denominator $\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})$ increases, resulting in a lower overall score. This log-based metric prevents inflated scores for candidates that might superficially fit many different instructions, thereby ensuring that each instruction is evaluated on a relatively fair basis. An ablation study (Appendix A.3) confirmed that omitting the logarithmic transformation significantly degrades normalization and impairs CodeRSA’s reranking capability, with Appendix A.4 offering geometric mean–based theoretical support.
Overall, at the pragmatic speaker level, CodeRSA leverages the code candidates themselves to sample potential instructions, capturing the diversity of possible interpretations. It then uses the literal listener’s estimation to quantify how effectively each potential instruction reflects the meaning of a code candidate.
Pragmatic Listener. The pragmatic listener (denoted $L_{1}$) re-examines the original instruction $i_{0}$ across all candidates, completing the backward reasoning guided by the pragmatic speaker’s preferences. In prior work, Degen (2023) defined a pragmatic listener as:
$\text{P}_{L_{1}}(c\mid i)\propto\text{P}_{S_{1}}(i\mid c)\cdot\text{P}(c),$
Here, $\text{P}(c)$ denotes the prior probability of a given code candidate. We make the simplifying assumption that all candidates are equally likely a priori, which allows us to treat $\text{P}(c)$ as a constant and omit it from the calculations. Note that this uniform prior assumption may not hold in real-world scenarios, as some code might be inherently more common than others. We will discuss this assumption further in Section 6. Consequently, in CodeRSA, the pragmatic listener reranks candidates by their pragmatic speaker scores:
$\text{P}_{L_{1}}(c\mid i)\propto\text{R}_{S_{1}}(i\mid c),$
so that the candidate with the highest pragmatic speaker score for $i_{0}$ is selected as optimal.
From CodeRSA’s reasoning process, it follows that, for each candidate, the LLM is invoked $m+1$ times, where $m$ represents the number of generated instructions. Consequently, the overall computational complexity of CodeRSA is $\mathcal{O}\bigl{(}n(m+1)\bigr{)}$, which grows quadratically when both $n$ and $m$ increase equally.
4 Experiment Setup
To analyze the merits of CodeRSA, we evaluate the performance of the three reranking methods (Coder, CoderReviewer, and CodeRSA) on a standard program code generation dataset using one of the latest language models. Since the strength of content-driven methods is their generality, we skipped extensive parameter tuning and simply used the default or commonly used parameters throughout our experiments.
4.1 Dataset and Model
The manually crafted HumanEval evaluation dataset (Chen et al., 2021) was designed to assess a model’s ability to convert natural language instructions into program code. It comprises 164 questions, each framed as an unfinished Python function starting with a brief instruction that describes how the function’s code should be completed.
We selected HumanEval for its proven role in code generation evaluation and balanced difficulty. For example, simpler datasets like CoNaLa (Yin et al., 2018) already yield near-perfect performance for the Coder model, leaving little room for reranking improvements. In contrast, more challenging datasets such as Bigcodebench (Zhuo et al., 2024) often produce many unsolvable instances, suggesting that more powerful models or improved problem modeling is needed.
Despite the availability of HumanEval for code generation, no dedicated and reliable benchmark currently exists for reranking generated code—a specialized subtask within code generation. Accordingly, the benchmark setup in our experiments involves three key stages: program code generation, testing (only for evaluation purposes), and reranking. Because LLM-based code generation is highly sensitive to hyperparameters and hardware configurations, we adopt a multi-sampling strategy. Specifically, for each of the 164 HumanEval problems, we first generate 300 candidate solutions at a temperature of 0.7. Then, we randomly select 50 problems, sample 10 candidates per problem, and repeat this process 10 times. Finally, we evaluate three different reranking methods on each resulting subset to measure their effectiveness.
In our experiments, we focus on the Llama-3-8B-Instruct variant, which comprises 8 billion parameters and is fine-tuned for instruction-following tasks Grattafiori et al. (2024). Llama-3-8B-Instruct is renowned for its lightweight design and robust performance, effectively balancing computational requirements with high-quality generation. Moreover, the HumanEval dataset poses a moderate challenge to the model, ensuring that it is neither overwhelmed by difficulty nor under-challenged. These characteristics render the model particularly suitable for evaluating the effectiveness of reranking methods in practical scenarios.
Additionally, we evaluated Llama-2-70b-chat Touvron et al. (2023) on a subset of HumanEval; however, its limited performance in code and instruction generation (and related poor probability estimates) led to its exclusion.
4.2 Implementation of Reranking Methods
Baselines. The Coder reranking method provides a straightforward way to compare the probability of a code candidate $c$ given the original instruction $i$. Specifically, it concatenates the instruction and code candidate in order (see Fig. 2, part A), prompting the language model to output token probabilities for the candidate sequentially. The product of these token probabilities then yields the cumulative probability of the entire code snippet. As mentioned in Section 3, Coder reranking can also be considered a literal listener-level approximation to $\text{P}(c\mid i)$; therefore, we use it as a baseline.
Another baseline, “random”, is defined as the ratio of correct candidates to total candidates in each subset. By using the expected proportion of correct codes, this approach minimizes the impact of random seeds and naturally reflects the inherent difficulty of each sub-dataset.
State-of-the-art Method. Zhang et al. (2023) showed that CoderReviewer reranking (see Section 2 for details) outperforms Coder reranking and rivals execution-driven methods such as CodeT.
Figure 2: The prompts used to calculate Coder score and generate additional instructions.
In practice, we use the same prompt format as in Coder reranking to compute $\text{P}(c\mid i)$. To compute $\text{P}(i\mid c)$, the order of the instruction and the generated code snippet is reversed in the prompt (see Appendix A.5.2).
CodeRSA. To balance running time and limited computing resources, we restrict the RSA reranking procedure in our experiments to 10 candidates per question. We then apply a one-shot prompt (see Fig. 2, part B) to the LLM to generate one instruction for each candidate, yielding a set of potential instructions $\{i_{0},i_{1},\ldots,i_{10}\}$, where $i_{0}$ is the original instruction. Consistent with Coder reranking’s input prompt, we then compute the probability of each code candidate $c^{\prime}$ under each potential instruction $i^{\prime}$, forming an $11\times 10$ Coder score matrix. Then, we derive pragmatic speaker scores for each candidate with respect to the original instruction by contrasting how the candidate performs under $i_{0}$ versus under the other generated instructions. Finally, we use these pragmatic scores to rerank the candidates, selecting the one with the highest pragmatic speaker score for $i_{0}$.
5 Results
5.1 Quantitative Analysis
In this section, we quantitatively evaluate three reranking methods, including CodeRSA, in terms of accuracy using ten subsets sampled from the HumanEval dataset and their corresponding generated code candidates.
Figure 3: Box plot of accuracy for the reranking methods. The orange line indicates the mean accuracy.
Here, accuracy is defined as the fraction of candidates selected by a reranking method that pass all test cases relative to the total number of candidates in the subset. Fig. 3 presents box plots comparing the accuracy distributions of the three reranking methods and a random baseline. Variance is estimated from bootstrapping across ten subsets. CodeRSA attains the highest mean accuracy, with relatively narrow interquartile ranges indicating consistent performance. CoderReviewer shows moderate accuracy, with a mean exceeding Coder’s but still below CodeRSA. Coder reranking exhibits a relatively large variance, indicating that its performance is less stable and more prone to fluctuations, likely due to its tendency to favor overly generic solutions.
(a) Details of question and two generated examples
(b) Coder Score Comparison
(c) Pragmatic Score Comparison
Figure 4: Two code candidates along with relevant information
Coder’s lower performance suggests that relying solely on a literal listener perspective is insufficient for effective reranking and that CodeRSA’s more complex reasoning leads to superior performance, demonstrating the benefits of a comprehensive RSA modeling approach.
5.2 Qualitative Analysis
Although our experiments show that CodeRSA achieves stable performance, it relies on certain idealized assumptions and an abstract reasoning process. To provide a more intuitive perspective, we include a qualitative analysis that examines how CodeRSA aligns with core RSA intuitions, thereby enhancing reranking quality.
Zhang et al. (2023) observed that language models often generate “degenerate” programs—overly generic yet repetitive code candidates. In our analysis, we observe that Coder reranking indeed favors these degenerate outputs because it evaluates the cumulative token likelihood of a candidate $c$, given the original instruction $i_{0}$:
$\text{P}(c\mid i_{0})\;=\;\prod_{t=1}^{|c|}\text{P}_{\text{LLM}}\bigl{(}c^{(t)}\mid i_{0},c^{(<t)}\bigr{)},$
where $c^{(t)}$ denotes the token at position $t$, and $c^{(<t)}$ represents the sequence of the preceding tokens. Each factor $\text{P}_{\text{LLM}}(c^{(t)}\mid i_{0},c^{(<t)})$ is strictly less than 1; thus, longer sequences accumulate a lower overall probability, biasing Coder reranking towards shorter, potentially degenerate code. Nonetheless, differences in token-level probabilities can offset this bias: a longer but logically correct candidate may achieve a higher overall probability if each token is assigned a sufficiently high probability.
With recent advances in LLMs, generating degenerate programs has become less common, although incomplete functionality remains a concern. For example, in Fig. 4, the original instruction requires returning the greatest integer above zero whose frequency is, at least, its own value, or $-1$ if none exists. Nevertheless, code_09 omits the “greatest” integer requirement and the $-1$ fallback, making it shorter and more likely to be highly ranked by Coder.
Fig. 4 presents the Coder scores $\log\text{P}_{L_{0}}(c\mid i)$ for function_01 and function_09 under all potential instructions (including $i_{0}$). Additionally, code_01 achieves a log-probability of $-29.24$ for $i_{0}$, whereas code_09 reaches $-21.12$. Consequently, Coder reranking selects the incomplete code_09. The Reviewer component alone is insufficient to offset code_09’s higher Coder score; hence CoderReviewer also prefers code_09. The difference between Coder and Reviewer scores highlights a major flaw in the CoderReviewer method. Since these scores often differ significantly, treating them as equally important is not optimal. Additionally, trying to adjust their weights can add more complexity and uncertainty.
Fig. 4 reports the pragmatic speaker scores $\text{R}_{S_{1}}(i\mid c)$ for each instruction–code pair. Notably, code_01 receives a score of $-0.0458$ with $i_{0}$, whereas code_09 has $-0.0533$. Acting as a pragmatic listener, CodeRSA selects code_01, which more closely aligns with $i_{0}$ from a pragmatic speaker’s perspective.
Although the score calculation in our method uses log probabilities, two observations confirm that our method still achieves the core intuitions put forward by the RSA framework:
Firstly, as can be seen in Fig. 4, code_01 has a lower overall Coder score than code_09, likely because it is longer and more complicated. However, under $i_{0}$, code_01 obtains a higher Coder score relative to other possible instructions, which then translates into a higher pragmatic speaker score after normalization. This process shows that CodeRSA examines each candidate’s relative scores across all potential instructions to verify whether it truly ranks highly under the original instruction.
Secondly, as shown in Fig. 4, code_09 receives multiple high Coder scores, especially under the instruction generated by itself. From the literal listener’s perspective, the self-generated instruction more precisely describes code_09, raising its overall Coder scores. Crucially, because the speaker’s score is normalized by the sum of log probabilities across all instructions for a given candidate, a higher Coder score on the self-generated instruction will reduce the relative weight assigned to the original instruction. Consequently, the pragmatic speaker will prefer to use another instruction for code_09 instead of the user’s original instruction. Moreover, as can be seen in Fig. 4, the self-generated instruction for code_09 matches its details more closely.
6 Discussion
Our proposed CodeRSA approach contains a number of simplifications compared to the original RSA model, which has been developed for describing human-human communication: (1) we assumed a uniform speaker cost for the instructions. While this simplification makes the analysis more tractable, it means that our model does not currently take into account effects related to how “costly” an instruction would be to produce for the human speaker. Future work should investigate variable cost structures to better capture these nuances.
We also assume that all code candidates are equally likely a priori, effectively omitting the candidate prior $\text{P}(c)$ from our calculations. Although this assumption allows us to concentrate solely on the pragmatic speaker’s score, it might not reflect real-world scenarios where some code candidates are more common due to usage frequency, domain-specific patterns, or contextual relevance. Future work should explore the use of non-uniform priors to better capture candidate plausibility.
In section 4, we argued that our approach, as a reranking approach, is most beneficial in situations where the dataset is not too easy (when a simple Coder model already achieves ceiling performance) and not too difficult, such that we can still obtain a high quality probability distribution over instructions and over code candidates. This raises the question of the relevance of pragmatic reasoning for code generation, and more generally in communication. Research on human communication has demonstrated the importance of pragmatic reasoning in communication despite its apparent computational overhead, while at the same time suggesting that humans may also learn to use simple heuristics or amortized estimates (Pu et al., 2024) to not engage in iterative reasoning in easy cases, while still being able to employ the full reasoning procedure in more difficult cases where it is necessary.
7 Conclusion
This work introduces CodeRSA, a candidate reranking algorithm for the generation of program code grounded in the Rational Speech Act framework. By modeling the iterative reasoning of a pragmatic listener about a pragmatic speaker, CodeRSA consistently outperforms the Coder reranking baseline and surpasses the state-of-the-art CoderReviewer approach. A qualitative analysis further reveals that, even when incorporating certain idealized assumptions and variations, CodeRSA remains faithful to the core principles of the RSA framework. These results highlight the effectiveness of applying well-established linguistic frameworks to enhance reasoning in language models, opening new avenues for research and development in code-related tasks.
8 Limitations
A known limitation of RSA approaches is their computational complexity and associated resource consumption. For example, on a single NVIDIA Tesla A100 (PCIe 4.0, 80GB HBM2e, 300W), performing complete CodeRSA inference on 500 instances takes nearly 8 hours. Our approach compares each potential instruction with every candidate, leading to a quadratic increase in complexity as the number of candidates grows. Although CodeRSA can theoretically handle many candidates, we limited our experiments to ten candidates per question to keep runtime and hardware usage manageable. This restriction inevitably narrows the variety of solutions and may affect how well the approach generalizes to larger-scale scenarios.
Reducing the computational overhead is a major goal for our future work. One promising direction is to design more lightweight scoring mechanisms or to adopt a multi-stage pipeline. For instance, a coarse filtering step could quickly discard low-probability solutions before applying CodeRSA’s full RSA-based reasoning to a smaller top-ranked subset. Alternatively, approximate models could reduce the number of token-level evaluations required, thereby preserving much of CodeRSA’s pragmatic reasoning benefits at a fraction of the computational cost. Such improvements would allow CodeRSA to scale more effectively and broaden its applicability to larger code generation tasks.
Another limitation is that while we employed a multi-sampling strategy to mitigate uncertainty, we have only utilized a single dataset and one model so far. We are currently working on incorporating several balanced-difficulty datasets—such as MBPP (Austin et al., 2021) and DS-1000 (Lai et al., 2023)—along with additional open-source models like Mistral (Jiang et al., 2023) and Qwen (Bai et al., 2023). This expansion will allow us to further assess the effectiveness of different reranking methods across diverse scenarios, ultimately leading to a more robust evaluation of our approach.
Appendix A Appendix
A.1 A Conjecture: Explaining CoderReviewer from an RSA Perspective
In the RSA framework, a pragmatic listener’s posterior over a candidate $c$ given an instruction $i$ is commonly expressed as:
$\text{P}_{L_{1}}(c\mid i)\propto\text{P}_{S_{1}}(i\mid c)\cdot\text{P}(c),$
where $\text{P}_{S_{1}}(i\mid c)$ represents how likely a pragmatic speaker would be to produce instruction $i$ when the correct candidate is $c$, and $\text{P}(c)$ is the prior likelihood of $c$.
Translating this perspective to LLMs, we hypothesize that when generating instructions (the “Reviewer” role), it is relatively straightforward for the model to produce abstract instructions from concrete code. Since code is unambiguous, the LLM can approximate a pragmatic speaker:
$\text{P}_{\text{LLM}}(i\mid c)\approx\text{P}_{S_{1}}(i\mid c).$
However, generating code from abstract instructions (the “Coder” role) is substantially more difficult. In this setting, the LLM may effectively revert to estimating a prior over possible candidates, thereby approximating:
$\text{P}_{\text{LLM}}(c\mid i)\approx\text{P}(c).$
From this RSA standpoint, the CoderReviewer paradigm can be considered a simplified, yet broad, modeling of a pragmatic listener.
A.2 More Details of Results
Table 1 summarizes the average accuracy and standard deviation across ten sampled subsets.
Table 1: Mean accuracy (with standard deviation) across 10 subsamples for each reranking method.
Method Mean Accuracy Std. Dev.
Coder 0.48 0.059
CoderReviewer 0.51 0.043
CodeRSA 0.55 0.044
random 0.43 0.030
As shown in Table 1, CodeRSA achieves the highest average accuracy (0.55), outperforming Coder (0.48), CoderReviewer (0.51), and the random baseline (0.43). This indicates that CodeRSA provides stronger overall performance compared to the other methods. From a standard-deviation perspective, CoderReviewer and CodeRSA both exhibit relatively stable performance across different subsamples. By contrast, Coder has the highest standard deviation (0.059), suggesting greater sensitivity to varying combinations of coding problems.
Figure 5: Performance difference between CodeRSA and CoderReviewer.
Fig. 5 depicts the per-trial accuracy difference between CodeRSA and CoderReviewer (i.e., CodeRSA accuracy minus CoderReviewer accuracy). Bars above zero indicate that CodeRSA outperforms CoderReviewer, while the single negative bar (Trial 2) represents the only instance where CodeRSA yields a lower accuracy. In most trials, CodeRSA demonstrates gains of up to 10%, highlighting its consistent advantage.
A.3 Ablation Study
Figure 6: Box plot of accuracy in ablation study.
To investigate the impact of the log transformation in the calculation of the pragmatic speaker score, we conducted an ablation study comparing the CodeRSA’s modeling of pragmatic speaker score:
$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})}\;,$
$\text{R}_{S_{1}}(i\mid c)\;\in\;(-1,\,0).$
against a variant that directly utilizes the normalized probability:
$\text{P}_{S_{1}}(i\mid c)=\frac{\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})},$
referred to as CodeRSA_nolog.
Fig. 6 displays the box plots of accuracy for different methods on ten subsets. From the results, CodeRSA achieves the highest mean accuracy with a narrow interquartile range, indicating stable performance. In contrast, CodeRSA_nolog shows a marked decrease in accuracy, evidenced by a lowest mean and a compressed interquartile range. This observation aligns with previous findings by Schuster et al. (2024), which reported that the basic RSA model underperforms the baseline on their SpreadNaLa dataset. These results suggest that applying a logarithmic transformation to expand the probability intervals can substantially enhance overall performance.
A.4 A Geometric Mean Perspective on the Log-Normalization Formula
In our proposed CodeRSA model, we define the pragmatic speaker score as follows:
$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})}\;,$
$\text{R}_{S_{1}}(i\mid c)\;\in\;(-1,\,0).$
where $\text{P}_{L_{0}}(c\mid i)$ is the probability estimated by the literal listener for candidate code $c$ given instruction $i$, and $I$ denotes the set of potential instructions. The geometric mean of the set $\{\text{P}_{L_{0}}(c\mid i^{\prime}):i^{\prime}\in I\}$ is defined as:
$\text{GM}=\left(\prod_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})\right)^{\frac{1}{|I|}},$
and its logarithm is given by:
$\log\text{GM}=\frac{1}{|I|}\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime}).$
Thus:
$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{|I|\cdot\log\text{GM}}\;.$
In this sense, normalizing the log probability for a particular instruction $i$ by the sum over all potential instructions can be seen as a way of capturing its relative contribution in a manner analogous to the geometric mean.
A.5 Prompt Used
A.5.1 For Generating the Additional Instruction:
##Write an instruction for given python function##
### Function start ###
def any_int(x, y, z):
if isinstance(x,int) and isinstance(y,int) and isinstance(z,int):
if (x+y==z) or (x+z==y) or (y+z==x):
return True
return False
return False
### Function end ###
### instruction start ###
Create a function that takes 3 numbers. Returns true if one of the numbers is equal to the sum of the other two, and all numbers are integers. Returns false in any other cases.
### instruction end ###
### Function start ###
any function
### Function end ###
###instruction start###
A.5.2 For Calculating the Reviewer Score (An Example):
def any_int(x, y, z):
if isinstance(x,int) and isinstance(y,int) and isinstance(z,int):
if (x+y==z) or (x+z==y) or (y+z==x):
return True
return False
return False
# Write a docstring for the above function
Create a function that takes 3 numbers. Returns true if one of the numbers is equal to the sum of the other two, and all numbers are integers. Returns false in any other cases.

### v3

@@
 However, user instructions often contain inherent ambiguities, making it challenging for LLMs to generate code that accurately reflects the user’s true intent.
-To address this challenge, researchers have proposed to produce multiple candidates of the program code and then rerank them to identify the best solution.
+To address this challenge, researchers have proposed approaches that produce multiple candidates of the program code and then rerank them to identify the best solution.
 In this paper, we propose CodeRSA, a novel code candidate reranking mechanism built upon the Rational Speech Act (RSA) framework, designed to guide LLMs toward more comprehensive pragmatic reasoning about user intent.
-We evaluate CodeRSA using one of the latest LLMs on a popular code generation dataset.
+We evaluate CodeRSA using Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct on two widely used code generation benchmarks, HumanEval and MBPP.
 Our experiment results show that CodeRSA consistently outperforms common baselines, surpasses the state-of-the-art approach in most cases, and demonstrates robust overall performance.
@@
 However, given the intrinsic complexities of coding and the potential ambiguities in user input, producing code in a single attempt may fail to explore the vast solution space, overlooking correct or higher-quality solutions (Liu et al., 2024a).
-A standard practice to address this shortcoming is to sample multiple solutions, which we refer to as “code candidates” (Chen et al., 2021; Brown et al., 2024), and to rerank them.
+A standard practice to address this shortcoming is to sample multiple solutions, which we refer to as code candidates (Chen et al., 2021; Brown et al., 2024), and to rerank them.
 Researchers have proposed various reranking strategies for code candidates, broadly divided into execution-driven and content-driven approaches.
-Due to the safety-risks associated with execution-driven approaches (Yetiştiren et al., 2023), we here focus on content-driven methods.
-These evaluate the generated text, often relying on token-level probabilities.
-For example, Coder reranking scores each candidate based on the cumulative probability of its tokens, sometimes however favoring “degenerate solutions” (generic or repetitive code) with disproportionately high token probabilities (Zhang et al., 2023).
-Figure 1: A comparison of our approach CodeRSA (top) compared to the CoderReviewer method (bottom).
+Due to the unreliability of automatically generated test suites (Chen et al., 2022) and the potential safety risks associated with code execution (Yetiştiren et al., 2023), we focus instead on content-driven methods, which evaluate the generated text, often relying on token-level probabilities.
+For example, Coder Reranking scores each candidate based on the cumulative probability of its tokens, sometimes however favoring ‘degenerate solutions” (generic or repetitive code) with disproportionately high token probabilities (Zhang et al., 2023a).
+Figure 1: A comparison of our approach CodeRSA (top) compared to CoderReviewer (bottom).
 When viewing code generation as a communicative process in which an LLM listens to the user’s intentions (Ouyang et al., 2022), Coder Reranking evaluates candidate solutions solely from the listener’s perspective.
 Yet, research on human communication suggests that effective listeners reason about the speaker (who in turn reasons about the listener) Grice (1975).
-Frank and Goodman (2012; 2016) provided a principled method for quantifying this process based on a probabilistic framework based on game-theoretic notions, called the Rational Speech Act (RSA) framework.
+Frank and Goodman (2012; 2016) provided a principled method for quantifying this process based on a probabilistic framework rooted in game-theoretic notions, called the Rational Speech Act (RSA) framework.
 Pu et al.
-(2020, 2024) demonstrated the RSA framework’s effectiveness on program generation for a simple domain, while Schuster et al.
+(2020, 2024) demonstrated the effectiveness of the RSA framework for program generation in a simple domain (regular-expression synthesis), while Schuster et al.
 (2024) reported negative results on a spreadsheet domain.
-One aspect that has held back RSA models from scaling up to realistic use cases is the computational overhead (Pu et al., 2024): it requires reasoning about the set of alternative instructions that the speaker could have given and about the set of alternative pieces of code that could solve the problem, which is very computationally expensive.
+One aspect that has held back RSA models from scaling up to realistic use cases is the computational overhead (Pu et al., 2024): It requires reasoning about a whole set of alternative instructions that the speaker could have given and about the set of alternative pieces of code that could solve the problem, which is computationally expensive.
 Zhang et al.
-(2023) therefore proposed CoderReviewer reranking as a simplified scalable approach that simplifies these probability estimation processes over alternatives.
-However, it comes at the cost of not fully modelling the dialogic, interactive reasoning that can emerge when speaker and listener exchange information.
-This work proposes CodeRSA, enabling LLMs to reason as pragmatic listeners and rank code candidates based on the user’s underlying intentions.
+(2023a) therefore proposed CoderReviewer Reranking as a scalable approach that simplifies these probability estimation processes over alternatives.
+However, this comes at the cost of not fully modelling the dialogic, interactive reasoning that can emerge when speaker and listener exchange information.
+In this paper, we propose CodeRSA, enabling LLMs to reason as pragmatic listeners and rank code candidates based on the user’s underlying intentions.
 It addresses the probability estimates for the set of alternative code candidates and alternative utterances via a sampling approach.
 CodeRSA generates multiple code candidates, and then generates additional instructions for each candidate, forming a set of potential instructions (including the original one), as illustrated in Fig.
-1: the literal listener $L_{0}$ estimates the probability of each code candidate given each instruction from the potential instruction set.
-The pragmatic speaker $S_{1}$ then applies a special normalization on these probabilities to quantify how specifically an instruction fits the generated code.
-By comparing these pragmatic speaker scores for the original instruction across all candidates, the pragmatic listener identifies the code candidate that most closely matches the original instruction from the speaker’s perspective, finalizing the reranking process (see Fig.
+1.
+Following the RSA framework, the literal listener $L_{0}$ first estimates the probability of each code candidate given each potential instruction.
+The pragmatic speaker $S_{1}$ then normalizes these probabilities to measure how specifically an instruction fits the generated code.
+Finally, by comparing these pragmatic speaker scores for the original instruction across all candidates, the pragmatic listener identifies the code candidate that best aligns with the user’s intent, completing the reranking process (see Fig.
 1).
-We conducted experiments using CodeRSA with Llama-3-8B-Instruct, one of the latest language models from the Llama family (Grattafiori et al., 2024) on OpenAI’s HumanEval benchmark (Chen et al., 2021).
-Our experiment results reveal that CodeRSA reliably outperforms the Coder and CoderReviewer reranking methods.
-Our qualitative analysis illustrates how the CodeRSA reranking enables better candidate selection, promoting a more comprehensive understanding of user intent.
+A challenge arises when many instructions are semantically equivalent but differ only in surface form.
+Applying RSA directly in such cases can lead to an overinterpretation of the formulation choice: The reasoning process is forced to treat near-identical descriptions as distinct alternatives, which were chosen for a reason of differentiating from other meaning alternatives.
+This fragments probability mass and reduces accuracy.
+To mitigate this, CodeRSA employs a clustering step.
+It groups semantically equivalent descriptions using an LLM-based equivalence test, ensuring that pragmatic reasoning emphasizes genuine differences in meaning rather than superficial wording (see Fig.
+1, Pragmatic Speaker Part, where the main cluster refers to the one containing the original description $I_{0}$).
+We conducted experiments using CodeRSA with Llama-3-8B-Instruct, one of the latest language models from the Llama family (Grattafiori et al., 2024), and Qwen-2.5-7B-Instruct, a recent instruction-tuned model from the Qwen series (Yang et al., 2024), on two widely used code generation benchmarks: OpenAI’s HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021).
+Our experimental results demonstrate that CodeRSA reliably outperforms the Coder and CoderReviewer Reranking methods.
+Our qualitative analysis reveals how the CodeRSA enables better candidate selection, promoting a more comprehensive understanding of user intent.
 2 Related Work
@@
 Their performance on code generation tasks often surpasses that of traditional models, and in many cases even rivals human programmers (Ni et al., 2024; Becker et al., 2023).
-Moreover, a recent study shows that LLMs also exhibit strong performance in code summarization, effectively translating code snippets into text (Akib et al., 2024).
-Diversity Sampling in Code Generation.
-In a prior study, Chen et al.
-(2021) found that allowing the model to generate more candidates significantly raises the probability of yielding a correct answer.
-This practice, which encourages the model to produce a broader range of potential outputs, is often referred to as diversity sampling.
+A recent study shows that LLMs also exhibit strong performance in code summarization, effectively translating code snippets into text (Akib et al., 2024).
 Code Reranking Methods.
-When LLMs generate multiple code candidates in response to the user instruction, execution-driven reranking methods such as CodeT (Chen et al., 2022) and AgentCoder (Huang et al., 2024) involve running test cases to gauge each candidate’s correctness.
-Although often effective, execution-driven approach can introduce additional safety risks and practical obstacles, such as the potential for malicious code execution, environment setup overhead, and resource constraints (Yetiştiren et al., 2023; Khoury et al., 2023).
+Execution-driven reranking methods such as CodeT (Chen et al., 2022) and AgentCoder (Huang et al., 2024) evaluate code candidates by running them against automatically generated test suites.
+Although often effective, these methods rely on the availability and reliability of test suites, which are frequently incomplete or difficult to construct, and executing untrusted code can pose safety risks (Yetiştiren et al., 2023; Khoury et al., 2023).
 In contrast, content-driven reranking methods are far more versatile because they do not rely on execution and are not even confined to coding tasks.
 Coder Reranking.
-In prior work, Chen et al.
-(2021) reranked code candidates by estimating $\text{P}(c\mid i)$, where $c$ denotes the generated code candidate and $i$ denotes the given instruction.
+Chen et al.
+(2021) rerank code candidates by estimating $\text{P}(c\mid i)$, where $c$ denotes the generated code candidate and $i$ denotes the given instruction.
 This process can also be called Coder Reranking because the LLM is a mere Coder that estimates the candidate probability based on the corresponding instruction.
 When using an LLM to estimate conditional probabilities, we compute the probability of each token iteratively.
-For example, in Coder reranking, the model processes candidate’s tokens from left to right: at each step, it calculates the probability of the current token given the instruction and the previously generated tokens, then appends that token to the context before moving on.
+For example, in Coder Reranking, the model processes a candidate’s tokens from left to right: At each step, it calculates the probability of the current token given the instruction and the previously generated tokens, then appends that token to the context before moving on.
 The product of these sequential probabilities across all tokens yields the overall probability of the code candidate under the given instruction:
@@
 Zhang et al.
-(2023) added the concept of a reviewer to Coder reranking (i.e., reassessing whether the instruction matches the generated code candidate).
-Formally, CoderReviewer metric is represented as follows:
-CoderReviewer $=\text{P}(c\mid i)\cdot\text{P}(i\mid c)$
+(2023a) introduced the idea of augmenting Coder Reranking with a reviewer, which jointly considers how likely a code candidate is under the instruction and how well the instruction is supported by the code.
+Formally, the CoderReviewer conditional probability is defined as:
+$\text{P}_{\text{CR}}(c\mid i)~\propto~\text{P}_{\text{LLM}}(c\mid i)\cdot\text{P}_{\text{LLM}}(i\mid c)$
 (Coder)    (Reviewer)
-The prompt positions of the instruction and code candidates are reversed in the reviewer’s implementation.
-Thus, the code-generation task is reformulated as an instruction-generation task to calculate the probability of a given instruction.
-The CoderReviewer metric is also considered a specialized form of maximum mutual information (Li and Jurafsky, 2016), measuring the bidirectional alignment between the generated code and the input instruction.
+By switching the positions of the instruction and code in the conditional formulation, the second term can be interpreted as reformulating the code-generation task as an instruction-generation task.
+This bidirectional formulation can be viewed as a specialized form of maximum mutual information (Li and Jurafsky, 2016).
 3 CodeRSA
-This section introduces CodeRSA, an approach that employs the Rational Speech Act (RSA) framework to enhance the reranking of candidate code snippets.
+In this section, we introduce CodeRSA, an approach that builds on the Rational Speech Act (RSA) framework to enhance the reranking of candidate code snippets.
 CodeRSA extends the models proposed by Cohn-Gordon et al.
@@
 The core innovation in CodeRSA arises from the pragmatic listener, which is responsible for selecting and reranking code candidates.
-It does so by imagining how a “pragmatic speaker” would choose an instruction that best distinguishes the intended code among various potential instructions.
-A pragmatic speaker evaluates how effectively a potential instruction communicates the intended code candidate.
-Within the RSA framework, this evaluation process necessitates an explicit calculation of probabilities.
-However, given the infinite combinations of possible instructions and code candidates, developing a robust and comprehensive probability estimation strategy becomes essential.
-To address this challenge, CodeRSA generates additional instructions for each candidate, thereby constructing an expanded set of potential instructions (including the original).
-Finally, by considering the potential instruction set, a pragmatic speaker can select the most appropriate instruction for each code candidate based on the motivation to effectively convey the message to a literal listener.
-At the foundation of CodeRSA is the literal listener, who estimates the probability of each code candidate by interpreting the user’s instruction word for word, without inferring deeper speaker intent.
+It does so by imagining how a pragmatic speaker would choose an instruction that best distinguishes the intended code among various potential instructions.
 Literal Listener.
 A literal listener (denoted $L_{0}$) represents the simplest level of reasoning in the RSA framework.
-It interprets utterances solely according to their literal meaning, without any higher-level pragmatic reasoning.
-Theoretically, let $c$ be a candidate code and $i$ be an instruction, then:
-$\text{P}_{L_{0}}(c\mid i)\propto\llbracket i\rrbracket(c),$
-where $\text{P}_{L_{0}}(c\mid i)$ denotes the literal listener’s estimation of probability of candidate $c$ given instruction $i$.
-The interpretation function $\llbracket i\rrbracket$ maps an instruction $i$ to a function that takes a code candidate $c$ and returns 1 if $c$ correctly implements $i$, and 0 otherwise.
-Likewise, most LLMs generate code by conditioning only on the user’s input instruction, which motivates CodeRSA’s use of an LLM to instantiate this literal listener:
+It interprets utterances solely according to their literal meaning, without any higher-level pragmatic inference.
+Let $c$ denote a candidate program and $i$ a user instruction.
+Then:
 $\text{P}_{L_{0}}(c\mid i)=\text{P}_{\text{LLM}}(c\mid i),$
-The literal listener provides CodeRSA’s baseline interpretation of user instructions, serving as the foundation for all subsequent reasoning.
-Note that the Coder reranking can also be considered a literal listener.
-For convenience, we refer to the logarithm of $\text{P}_{L_{0}}(c\mid i)$ as the Coder score.
+where $\text{P}_{\text{LLM}}(c\mid i)$ is the probability assigned by the LLM to candidate $c$ given instruction $i$.
+In an idealized RSA setting, the literal listener would evaluate all possible programs, but since the space of programs is unbounded, we approximate it by sampling a finite set of candidate codes from the LLM.
+We additionally define a candidate prior distribution obtained by querying the LLM without any instruction context:
+$\text{P}_{\text{prior}}(c)=\text{P}_{\text{LLM}}(c\mid\varnothing).$
+This prior reflects how plausible a candidate program is in general, independent of the specific user instruction.
 Pragmatic Speaker.
-In the RSA framework, the pragmatic speaker (denoted $S_{1}$) is primarily responsible for determining whether an instruction $i$ effectively conveys the intended meaning of a candidate code $c$ to the literal listener.
+In the RSA framework, the pragmatic speaker (denoted $S_{1}$) is primarily responsible for determining whether an instruction $i$ effectively conveys the intended meaning of a candidate $c$ to the literal listener.
 Formally, a pragmatic speaker can be defined as:
-$\text{P}_{S_{1}}(i\mid c)=\frac{\exp\left(\log\text{P}_{L_{0}}(c\mid i)-C(i)\right)}{\sum_{i^{\prime}}\exp\left(\log\text{P}_{L_{0}}\left(c\mid i^{\prime}\right)-C\left(i^{\prime}\right)\right)},$
-Here, $\operatorname{C}(i)$ denotes a cost function that quantifies the “expense” of employing a particular instruction.
-This formula quantifies the pragmatic speaker’s estimation of the probability that a particular instruction will be used to describe a code candidate, as interpreted from the literal listener’s perspective.
-Note that an idealized RSA approach would consider every possible instruction $i^{\prime}$, which is infeasible in practice.
-Instead, CodeRSA leverages the fact that ambiguities in the original instruction often surface in the code candidates generated by LLMs, thereby enabling a sample-based estimation.
-Given a user-provided instruction $i_{0}$, we request multiple candidate solutions $\{c_{1},\dots,c_{n}\}$ from the LLM.
-Each candidate $c_{j}$ may capture a slightly different interpretation of $i_{0}$.
-We then derive one or more instructions from each candidate $c_{j}$, forming a set of potential instructions $\{i_{0},i_{1},\ldots,i_{m}\}$, which we denote by $I$.
-This approach ensures that we sample relevant alternative instructions directly tied to how the model interprets the original instruction.
-Potential instructions then constitute a finite sample set that approximates the otherwise infinite instruction space for pragmatic speaker estimation.
+$\text{P}_{S_{1}}(i\mid c)=\frac{\exp\left(\log\text{P}_{L_{0}}(c\mid i)-C(i)\right)}{\sum_{i^{\prime}}\exp\left(\log\text{P}_{L_{0}}\left(c\mid i^{\prime}\right)-C\left(i^{\prime}\right)\right)}.$
+Here, $C(i)$ denotes a cost function for using instruction $i$.
+In an ideal RSA setting, the normalization spans every possible instruction $i^{\prime}$, which is intractable for code generation.
+To approximate this space in practice, we take the sampled candidate codes as anchors and derive $m$ alternative instructions from each of the $n$ code candidates, together with the original instruction $i_{0}$, yielding a finite task-relevant instruction set $I=\{i_{0},i_{1},\ldots,i_{mn}\}$.
+This construction provides a principled approximation of the otherwise infinite instruction space while keeping RSA’s normalization meaningful.
 To simplify the model and focus on core pragmatic reasoning, we assume a uniform cost for all instructions, which effectively cancels out during normalization.
-However, a more detailed modeling of the cost function may provide additional insights, a point we further discuss in the Section 6.
+A detailed modeling of the cost function may provide additional insights, a point we further discuss in Section 6.
 A pragmatic speaker then can be defined in a simplified form as:
-$\text{P}_{S_{1}}(i\mid c)=\frac{\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})},$
-In practice, the distribution of $\text{P}_{L_{0}}(c\mid i)$ is often highly skewed, with a few high probability candidates dominating and a large spread in probability estimates.
-To address this, we here propose to instead work with log probabilities, which in practice brings low probability alternatives closer together and gives more weight to the relative trends among instruction–code pairs.
-The pragmatic speaker’s preference score for a given instruction is hence quantified as:
-$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})}\;,$
-$\text{R}_{S_{1}}(i\mid c)\;\in\;(-1,\,0).$
-In log space, after normalizing with respect to log probabilities, smallest numbers will now correspond to the events with the highest probability.
-Therefore, we invert the sign to negative in order to still choose maximal score instructions.
-Notably, if $c$ is overly generic or appears plausible under multiple instructions, then the denominator $\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})$ increases, resulting in a lower overall score.
-This log-based metric prevents inflated scores for candidates that might superficially fit many different instructions, thereby ensuring that each instruction is evaluated on a relatively fair basis.
-An ablation study (Appendix A.3) confirmed that omitting the logarithmic transformation significantly degrades normalization and impairs CodeRSA’s reranking capability, with Appendix A.4 offering geometric mean–based theoretical support.
-Overall, at the pragmatic speaker level, CodeRSA leverages the code candidates themselves to sample potential instructions, capturing the diversity of possible interpretations.
-It then uses the literal listener’s estimation to quantify how effectively each potential instruction reflects the meaning of a code candidate.
+$\text{P}_{S_{1}}(i\mid c)=\frac{\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})}.$
 Pragmatic Listener.
 The pragmatic listener (denoted $L_{1}$) re-examines the original instruction $i_{0}$ across all candidates, completing the backward reasoning guided by the pragmatic speaker’s preferences.
-In prior work, Degen (2023) defined a pragmatic listener as:
+In the standard RSA formulation (Degen, 2023), a pragmatic listener is defined as:
 $\text{P}_{L_{1}}(c\mid i)\;\propto\;\text{P}_{S_{1}}(i\mid c)\cdot\text{P}(c),$
-Here, $\text{P}(c)$ denotes the prior probability of a given code candidate.
-We make the simplifying assumption that all candidates are equally likely a priori, which allows us to treat $\text{P}(c)$ as a constant and omit it from the calculations.
-Note that this uniform prior assumption may not hold in real-world scenarios, as some code might be inherently more common than others.
-We will discuss this assumption further in Section 6.
-Consequently, in CodeRSA, the pragmatic listener reranks candidates by their pragmatic speaker scores:
-$\text{P}_{L_{1}}(c\mid i)\propto\text{R}_{S_{1}}(i\mid c),$
-so that the candidate with the highest pragmatic speaker score for $i_{0}$ is selected as optimal.
-From CodeRSA’s reasoning process, it follows that, for each candidate, the LLM is invoked $m+1$ times, where $m$ represents the number of generated instructions.
-Consequently, the overall computational complexity of CodeRSA is $\mathcal{O}\bigl{(}n(m+1)\bigr{)}$, which grows quadratically when both $n$ and $m$ increase equally.
+where $\text{P}(c)$ denotes the prior probability of candidate $c$.
+In practice, directly multiplying a normalized distribution by $\text{P}(c)$ can distort the allocation of probability mass, as the prior may dominate post hoc.
+Instead, CodeRSA incorporates priors via a candidate-specific temperature applied before normalization at the speaker stage.
+Let $z_{c}$ be the within-task standardized log prior of candidate $c$ (estimated from the LLM without conditioning context), and define a candidate-specific temperature as
+$\tau_{c}\;=\;e^{-\alpha z_{c}},\qquad\alpha\geq 0,\quad\tau_{c}>0,$
+where $\alpha$ controls how strongly the prior influences the temperature scaling.
+A higher prior (larger $z_{c}$) yields a smaller temperature ($\tau_{c}<1$) and thus a sharper distribution over alternatives.
+Candidates with higher priors therefore emphasize their most confident clusters more strongly, typically those that align best with the original instruction (e.g., the “main cluster”), giving them a comparative advantage during reranking.
+With this calibration, the pragmatic speaker used by the listener is
+$\text{P}_{S_{1}}(i\mid c;\tau_{c})\;=\;\frac{\bigl(\text{P}_{L_{0}}(c\mid i)\bigr)^{1/\tau_{c}}}{\sum_{i^{\prime}\in I}\bigl(\text{P}_{L_{0}}(c\mid i^{\prime})\bigr)^{1/\tau_{c}}},$
+which reduces to the standard RSA speaker when $\alpha=0$ (thus $\tau_{c}=1$). Finally, the pragmatic listener ranks candidates with respect to the original instruction:
+$\text{P}_{L_{1}}(c\mid i_{0})\;\propto\;\text{P}_{S_{1}}(i_{0}\mid c;\tau_{c}).$
+This formulation preserves the spirit of RSA while integrating priors in a stable and interpretable manner: Rather than post-hoc reweighting, priors act as adaptive temperatures that shape the pragmatic reasoning process upstream of normalization.
+In our experiments, we treat $\alpha$ as a tunable hyperparameter and find that performance is stable across a broad range of values (see Section 5).
+Clustering Paraphrases.
+While the basic RSA formulation operates directly over the instruction set $I$, it can suffer from over-interpreting superficial variations in wording when many instructions are semantically equivalent and differ only in surface form.
+In such cases, RSA allocates probability mass across paraphrases as if they were meaningful distinct alternatives, diluting the signal and reducing accuracy.
+Semantically equivalent instructions:
+ “return the sum of a list of integers”
+ “compute the total of all integers in a list”
+Non-equivalent instruction:
+ “return the product of a list of integers”
+Table 1: Examples of equivalent and non-equivalent instructions from MBPP. The first group expresses the same semantics, while the second differs in meaning.
+To mitigate this, CodeRSA employs a semantic clustering.
+Candidate instructions are grouped into semantic clusters $\mathcal{C}=\{C_{1},\ldots,C_{K}\}$ using an LLM-based equivalence test (implementation details in Section 4), so that pragmatic reasoning operates over clusters rather than individual instructions.
+This ensures that comparisons emphasize genuine differences in meaning rather than superficial variation.
+For a candidate $c$ and cluster $C_{k}$, the literal listener probability is aggregated as:
+$\text{P}_{L_{0}}(c\mid C_{k})=\begin{cases}\text{P}_{L_{0}}(c\mid i_{0}),&\text{if }i_{0}\in C_{k},\\[4.0pt] \frac{1}{|C_{k}|}\!\sum_{i\in C_{k}}\!\text{P}_{L_{0}}(c\mid i),&\text{otherwise.}\end{cases}$
+The pragmatic speaker distribution over clusters then becomes:
+$\text{P}_{S_{1}}(C_{k}\mid c;\tau_{c})\;=\;\frac{\bigl(\text{P}_{L_{0}}(c\mid C_{k})\bigr)^{1/\tau_{c}}}{\sum_{C_{k^{\prime}}\in\mathcal{C}}\bigl(\text{P}_{L_{0}}(c\mid C_{k^{\prime}})\bigr)^{1/\tau_{c}}},$
+where the candidate-specific temperature $\tau_{c}=e^{-\alpha z_{c}}$ incorporates priors.
+Finally, the pragmatic listener reranks candidates with respect to the cluster $C^{\ast}$ containing the original instruction $i_{0}$:
+$\text{P}_{L_{1}}(c\mid i_{0})\;\propto\;\text{P}_{S_{1}}(C^{\ast}\mid c;\tau_{c}).$
+This extension preserves the primacy of the original instruction while preventing RSA from over-differentiating among paraphrases.
+Moreover, the integration of priors through adaptive temperatures ensures that the reranking remains calibrated against candidate plausibility.
 4 Experiment Setup
-To analyze the merits of CodeRSA, we evaluate the performance of the three reranking methods (Coder, CoderReviewer, and CodeRSA) on a standard program code generation dataset using one of the latest language models.
-Since the strength of content-driven methods is their generality, we skipped extensive parameter tuning and simply used the default or commonly used parameters throughout our experiments.
-4.1 Dataset and Model
-The manually crafted HumanEval evaluation dataset (Chen et al., 2021) was designed to assess a model’s ability to convert natural language instructions into program code.
-It comprises 164 questions, each framed as an unfinished Python function starting with a brief instruction that describes how the function’s code should be completed.
-We selected HumanEval for its proven role in code generation evaluation and balanced difficulty.
-For example, simpler datasets like CoNaLa (Yin et al., 2018) already yield near-perfect performance for the Coder model, leaving little room for reranking improvements.
-In contrast, more challenging datasets such as Bigcodebench (Zhuo et al., 2024) often produce many unsolvable instances, suggesting that more powerful models or improved problem modeling is needed.
-Despite the availability of HumanEval for code generation, no dedicated and reliable benchmark currently exists for reranking generated code—a specialized subtask within code generation.
-Accordingly, the benchmark setup in our experiments involves three key stages: program code generation, testing (only for evaluation purposes), and reranking.
-Because LLM-based code generation is highly sensitive to hyperparameters and hardware configurations, we adopt a multi-sampling strategy.
-Specifically, for each of the 164 HumanEval problems, we first generate 300 candidate solutions at a temperature of 0.7.
-Then, we randomly select 50 problems, sample 10 candidates per problem, and repeat this process 10 times.
-Finally, we evaluate three different reranking methods on each resulting subset to measure their effectiveness.
-In our experiments, we focus on the Llama-3-8B-Instruct variant, which comprises 8 billion parameters and is fine-tuned for instruction-following tasks Grattafiori et al.
-(2024).
-Llama-3-8B-Instruct is renowned for its lightweight design and robust performance, effectively balancing computational requirements with high-quality generation.
-Moreover, the HumanEval dataset poses a moderate challenge to the model, ensuring that it is neither overwhelmed by difficulty nor under-challenged.
-These characteristics render the model particularly suitable for evaluating the effectiveness of reranking methods in practical scenarios.
-Additionally, we evaluated Llama-2-70b-chat Touvron et al.
-(2023) on a subset of HumanEval; however, its limited performance in code and instruction generation (and related poor probability estimates) led to its exclusion.
+To understand the strengths and weaknesses of CodeRSA, we evaluate the performance of three reranking methods (Coder, CoderReviewer, and CodeRSA) on widely used benchmarks for code generation.
+Since the advantage of content-driven methods lies in their generality, we rely on commonly adopted default settings and perform only minimal sensitivity checks on key parameters.
+4.1 Dataset and Base Models
+We evaluate on two widely used code generation benchmarks.
+HumanEval (Chen et al., 2021) contains 164 Python programming problems, each presented as an unfinished function with a natural language instruction.
+MBPP (Austin et al., 2021) includes 257 short programming tasks with natural language prompts.
+HumanEval offers balanced difficulty, while MBPP introduces greater lexical variety.
+Note that simpler datasets such as CoNaLa (Yin et al., 2018) already yield near-perfect results, leaving little room for reranking, whereas more challenging datasets such as BigCodeBench (Zhuo et al., 2024) contain many instances that cannot yet be solved by today’s state-of-the-art models, which makes it difficult to obtain meaningful comparisons of reranking methods and may obscure the performance differences we aim to study.
+We use the following setup: for each problem in HumanEval and MBPP, we sample $n=10$ candidate codes at a temperature of 1.0.
+We then evaluate reranking methods on this shared candidate set.
+A sensitivity check with varying numbers of sampled candidates is provided in Appendix A.3.
+For our experiments, we use Llama-3-8B-Instruct (Grattafiori et al., 2024) and Qwen-2.5-7B-Instruct (Yang et al., 2024), two instruction-tuned LLMs of comparable scale.
+Llama-3-8B-Instruct balances efficiency with strong generation quality, while Qwen-2.5-7B-Instruct provides a competitive open-source alternative.
+Both achieve competitive performance on HumanEval and MBPP, making them suitable for assessing reranking in our setting.
+We do not include specialized coder models in this study, since our framework requires both code generation and instruction-level reasoning.
+Future work could explore hybrid setups, for example, using coder models for program synthesis combined with general-purpose instruction-tuned models for reasoning about instructions.
 4.2 Implementation of Reranking Methods
@@
 As mentioned in Section 3, Coder Reranking can also be considered a literal listener-level approximation to $\text{P}(c\mid i)$; therefore, we use it as a baseline.
-Another baseline, “random”, is defined as the ratio of correct candidates to total candidates in each subset.
-By using the expected proportion of correct codes, this approach minimizes the impact of random seeds and naturally reflects the inherent difficulty of each sub-dataset.
-State-of-the-art Method. Zhang et al. (2023) showed that CoderReviewer reranking (see Section 2 for details) outperforms Coder reranking and rivals execution-driven methods such as CodeT.
-Figure 2: The prompts used to calculate Coder score and generate additional instructions.
+State-of-the-art Method.
+Zhang et al.
+(2023a) showed that CoderReviewer Reranking (see Section 2 for details) outperforms Coder Reranking and rivals execution-driven methods such as CodeT.
 In practice, we use the same prompt format as in Coder Reranking to compute $\text{P}(c\mid i)$.
 To compute $\text{P}(i\mid c)$, the order of the instruction and the generated code snippet is reversed in the prompt (see Appendix A.5.2).
+Figure 2: The prompts used to calculate Coder score and generate additional instructions.
 CodeRSA.
-To balance running time and limited computing resources, we restrict the RSA reranking procedure in our experiments to 10 candidates per question.
-We then apply a one-shot prompt (see Fig.
-2, part B) to the LLM to generate one instruction for each candidate, yielding a set of potential instructions $\{i_{0},i_{1},\ldots,i_{10}\}$, where $i_{0}$ is the original instruction.
-Consistent with Coder reranking’s input prompt, we then compute the probability of each code candidate $c^{\prime}$ under each potential instruction $i^{\prime}$, forming an $11\times 10$ Coder score matrix.
-Then, we derive pragmatic speaker scores for each candidate with respect to the original instruction by contrasting how the candidate performs under $i_{0}$ versus under the other generated instructions.
-Finally, we use these pragmatic scores to rerank the candidates, selecting the one with the highest pragmatic speaker score for $i_{0}$.
+To balance runtime and computational constraints, we limit the process to $n=10$ candidate programs per problem.
+For CodeRSA, we further generate one additional instruction ($m=1$) for each candidate using a one-shot prompt (temperature = 0.7; see Fig.
+2, part B).
+Rather than treating each instruction independently, we next use the same LLM that generated the candidates to perform pairwise semantic equivalence judgments among these instructions.
+Following prior work on LLM-based clustering (Zhang et al., 2023b), we cluster instructions that express the same functionality (see Section 3).
+Concretely, we query the LLM with a 3-shot prompt containing examples of both positive and negative semantic equivalence pairs, asking it to judge whether two instructions express the same functionality.
+We then build a pairwise equivalence graph where each node represents an instruction and an edge indicates semantic equivalence according to the LLM.
+The connected components of this graph are treated as clusters of mutually equivalent instructions (see Appendix A.4).
+For each candidate $c$, we compute literal listener scores with respect to every instruction, and then aggregate them at the cluster level: non-main clusters take the mean across their members, while the cluster containing the original instruction $i_{0}$ retains its direct probability.
+We also incorporate candidate priors through a candidate-specific temperature parameter $\tau_{c}$, controlled by a coefficient $\alpha$ (see Section 3 for the definition).
+Finally, we apply softmax normalization with these temperatures over clusters to obtain cluster-level pragmatic speaker scores.
+The pragmatic listener then reranks candidates by selecting the one with the highest speaker score with respect to the $i_{0}$-cluster, which represents the original user intent.
 5 Results
 5.1 Quantitative Analysis
-In this section, we quantitatively evaluate three reranking methods, including CodeRSA, in terms of accuracy using ten subsets sampled from the HumanEval dataset and their corresponding generated code candidates.
-Figure 3: Box plot of accuracy for the reranking methods. The orange line indicates the mean accuracy.
-Here, accuracy is defined as the fraction of candidates selected by a reranking method that pass all test cases relative to the total number of candidates in the subset.
-Fig.
-3 presents box plots comparing the accuracy distributions of the three reranking methods and a random baseline.
-Variance is estimated from bootstrapping across ten subsets.
-CodeRSA attains the highest mean accuracy, with relatively narrow interquartile ranges indicating consistent performance.
-CoderReviewer shows moderate accuracy, with a mean exceeding Coder’s but still below CodeRSA.
-Coder reranking exhibits a relatively large variance, indicating that its performance is less stable and more prone to fluctuations, likely due to its tendency to favor overly generic solutions.
+In this section, we analyze the quantitative performance of CodeRSA with respect to the calibration parameter $\alpha$ (defined in Section 3), which controls the influence of the prior through temperature scaling, using the MBPP dataset and Llama-3-8B-Instruct model.
+Figure 3: Accuracy of CodeRSA across different values of the calibration parameter $\alpha$. The shaded region indicates a stable performance band.
+Here, accuracy denotes the proportion of test instances where the reranking method selects a candidate that passes all test cases provided by the benchmark.
+Figure 3 shows that CodeRSA consistently outperforms baseline reranking methods, with clustering playing a crucial role: removing the clustering step yields a substantial drop in accuracy.
+The figure also highlights the robustness of CodeRSA to the choice of $\alpha$: within the stable band of $[0.90,1.15]$, performance remains consistently above both Coder and CoderReviewer.
+At $\alpha=1.0$, CodeRSA achieves the best accuracy of 59.53%, clearly surpassing the baselines.
+These results demonstrate that CodeRSA’s pragmatic reasoning, enhanced by clustering, is not overly sensitive to calibration.
+We report further results covering different models (Qwen2.5-7B-Instruct) and datasets (HumanEval) in Appendix A.2, which confirm the same overall trends.
 (a) Details of question and two generated examples
@@
 (c) Pragmatic Score Comparison
-Figure 4: Two code candidates along with relevant information
-Coder’s lower performance suggests that relying solely on a literal listener perspective is insufficient for effective reranking and that CodeRSA’s more complex reasoning leads to superior performance, demonstrating the benefits of a comprehensive RSA modeling approach.
+Figure 4: Qualitative Example: Bias in Coder vs. CodeRSA Correction
 5.2 Qualitative Analysis
@@
 Zhang et al.
-(2023) observed that language models often generate “degenerate” programs—overly generic yet repetitive code candidates.
-In our analysis, we observe that Coder reranking indeed favors these degenerate outputs because it evaluates the cumulative token likelihood of a candidate $c$, given the original instruction $i_{0}$:
-$\text{P}(c\mid i_{0})\;=\;\prod_{t=1}^{|c|}\text{P}_{\text{LLM}}\bigl{(}c^{(t)}\mid i_{0},c^{(<t)}\bigr{)},$
-where $c^{(t)}$ denotes the token at position $t$, and $c^{(<t)}$ represents the sequence of the preceding tokens.
-Each factor $\text{P}_{\text{LLM}}(c^{(t)}\mid i_{0},c^{(<t)})$ is strictly less than 1; thus, longer sequences accumulate a lower overall probability, biasing Coder reranking towards shorter, potentially degenerate code.
-Nonetheless, differences in token-level probabilities can offset this bias: a longer but logically correct candidate may achieve a higher overall probability if each token is assigned a sufficiently high probability.
-With recent advances in LLMs, generating degenerate programs has become less common, although incomplete functionality remains a concern.
-For example, in Fig.
-4, the original instruction requires returning the greatest integer above zero whose frequency is, at least, its own value, or $-1$ if none exists.
-Nevertheless, code_09 omits the “greatest” integer requirement and the $-1$ fallback, making it shorter and more likely to be highly ranked by Coder.
+(2023a) note that reranking based on cumulative token likelihood tends to prefer shorter candidates, since each token probability is <1<1 and longer sequences accumulate lower overall scores.
+This bias makes the Coder approach prone to favoring incomplete or generic programs.
+In Fig.
+4, the instruction requires returning the greatest integer above zero whose frequency is at least its own value, or 1-1 if none exists.
+However, code_09 omits both the “greatest” requirement and the 1-1 fallback, making it incomplete but shorter.
+As shown in Fig.
+4, Coder assigns code_09 a higher score (21.12-21.12) than the correct code_01 (29.24-29.24), and thus prefers the degenerate solution.
+CoderReviewer inherits this issue, as Reviewer alone cannot offset code_09’s inflated Coder score.
 Fig.
-4 presents the Coder scores $\log\text{P}_{L_{0}}(c\mid i)$ for function_01 and function_09 under all potential instructions (including $i_{0}$).
-Additionally, code_01 achieves a log-probability of $-29.24$ for $i_{0}$, whereas code_09 reaches $-21.12$.
-Consequently, Coder reranking selects the incomplete code_09.
-The Reviewer component alone is insufficient to offset code_09’s higher Coder score; hence CoderReviewer also prefers code_09.
-The difference between Coder and Reviewer scores highlights a major flaw in the CoderReviewer method.
-Since these scores often differ significantly, treating them as equally important is not optimal.
-Additionally, trying to adjust their weights can add more complexity and uncertainty.
-Fig.
-4 reports the pragmatic speaker scores $\text{R}_{S_{1}}(i\mid c)$ for each instruction–code pair.
-Notably, code_01 receives a score of $-0.0458$ with $i_{0}$, whereas code_09 has $-0.0533$.
-Acting as a pragmatic listener, CodeRSA selects code_01, which more closely aligns with $i_{0}$ from a pragmatic speaker’s perspective.
-Although the score calculation in our method uses log probabilities, two observations confirm that our method still achieves the core intuitions put forward by the RSA framework:
-Firstly, as can be seen in Fig.
-4, code_01 has a lower overall Coder score than code_09, likely because it is longer and more complicated.
-However, under $i_{0}$, code_01 obtains a higher Coder score relative to other possible instructions, which then translates into a higher pragmatic speaker score after normalization.
-This process shows that CodeRSA examines each candidate’s relative scores across all potential instructions to verify whether it truly ranks highly under the original instruction.
-Secondly, as shown in Fig.
-4, code_09 receives multiple high Coder scores, especially under the instruction generated by itself.
-From the literal listener’s perspective, the self-generated instruction more precisely describes code_09, raising its overall Coder scores.
-Crucially, because the speaker’s score is normalized by the sum of log probabilities across all instructions for a given candidate, a higher Coder score on the self-generated instruction will reduce the relative weight assigned to the original instruction.
-Consequently, the pragmatic speaker will prefer to use another instruction for code_09 instead of the user’s original instruction.
-Moreover, as can be seen in Fig.
-4, the self-generated instruction for code_09 matches its details more closely.
+4 reports CodeRSA’s cluster-level pragmatic speaker scores after softmax normalization ($\alpha=1$).
+Instructions with equivalent semantics are grouped, and the cluster containing the original instruction $i_{0}$ is treated as the main cluster.
+Here, code_01 achieves a score of 0.32 on the main cluster, compared to 0.25 for code_09.
+Notably, code_09 also receives relatively high confidence on Cluster_4, which dilutes its probability on the main cluster due to RSA normalization.
+In RSA terms, code_09 is not strongly aligned with either the main cluster or Cluster_4, indicating that it fits the intended instruction less well than other candidates.
+By contrast, the probability of code_01 is concentrated on the main cluster, which better aligns with the original instruction and is therefore favored under pragmatic reasoning.
+Taken together, this case study shows how CodeRSA operationalizes RSA reasoning: By normalizing over alternative clusters, it penalizes candidates that spread probability mass across multiple interpretations and favors those that focus on the main cluster, thereby improving robustness and faithfulness in reranking.
 6 Discussion
-Our proposed CodeRSA approach contains a number of simplifications compared to the original RSA model, which has been developed for describing human-human communication: (1) we assumed a uniform speaker cost for the instructions.
+Our proposed CodeRSA approach contains a number of simplifications compared to the original RSA model, which has been developed for describing human–human communication: It assumes a uniform speaker cost for the instructions.
 While this simplification makes the analysis more tractable, it means that our model does not currently take into account effects related to how “costly” an instruction would be to produce for the human speaker.
 Future work should investigate variable cost structures to better capture these nuances.
-We also assume that all code candidates are equally likely a priori, effectively omitting the candidate prior $\text{P}(c)$ from our calculations.
-Although this assumption allows us to concentrate solely on the pragmatic speaker’s score, it might not reflect real-world scenarios where some code candidates are more common due to usage frequency, domain-specific patterns, or contextual relevance.
-Future work should explore the use of non-uniform priors to better capture candidate plausibility.
-In section 4, we argued that our approach, as a reranking approach, is most beneficial in situations where the dataset is not too easy (when a simple Coder model already achieves ceiling performance) and not too difficult, such that we can still obtain a high quality probability distribution over instructions and over code candidates.
+In Section 4, we argued that CodeRSA, as a reranking approach, is most beneficial in situations where the dataset is not too easy (when a simple Coder model already achieves ceiling performance) and not too difficult, such that we can still obtain a high quality probability distribution over instructions and over code candidates.
 This raises the question of the relevance of pragmatic reasoning for code generation, and more generally in communication.
-Research on human communication has demonstrated the importance of pragmatic reasoning in communication despite its apparent computational overhead, while at the same time suggesting that humans may also learn to use simple heuristics or amortized estimates (Pu et al., 2024) to not engage in iterative reasoning in easy cases, while still being able to employ the full reasoning procedure in more difficult cases where it is necessary.
+Research on human communication has demonstrated the importance of pragmatic reasoning, even though it introduces additional computational overhead.
+At the same time, studies suggest that humans may rely on simple heuristics or amortized estimates (Pu et al., 2024), avoiding iterative reasoning in easy cases while still engaging in full pragmatic reasoning when tasks are more complex.
 7 Conclusion
@@
 A known limitation of RSA approaches is their computational complexity and associated resource consumption.
-For example, on a single NVIDIA Tesla A100 (PCIe 4.0, 80GB HBM2e, 300W), performing complete CodeRSA inference on 500 instances takes nearly 8 hours.
+For example, on a single NVIDIA Tesla A100 (PCIe 4.0, 80GB HBM2e, 300W), performing complete CodeRSA inference on 500 instances takes nearly 6 hours.
 Our approach compares each potential instruction with every candidate, leading to a quadratic increase in complexity as the number of candidates grows.
@@
 Such improvements would allow CodeRSA to scale more effectively and broaden its applicability to larger code generation tasks.
-Another limitation is that while we employed a multi-sampling strategy to mitigate uncertainty, we have only utilized a single dataset and one model so far.
-We are currently working on incorporating several balanced-difficulty datasets—such as MBPP (Austin et al., 2021) and DS-1000 (Lai et al., 2023)—along with additional open-source models like Mistral (Jiang et al., 2023) and Qwen (Bai et al., 2023).
-This expansion will allow us to further assess the effectiveness of different reranking methods across diverse scenarios, ultimately leading to a more robust evaluation of our approach.
+Another limitation is that, although our experiments already cover two models (Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct) and two datasets (HumanEval and MBPP), the scope remains relatively narrow.
+We are currently working on incorporating additional balanced-difficulty datasets such as DS-1000 (Lai et al., 2023)), along with further open-source models like Mistral (Jiang et al., 2023) and newer Qwen releases beyond Qwen-2.5.
+This expansion will allow us to evaluate reranking methods across a wider range of scenarios, ultimately leading to a more comprehensive assessment of our approach.
 Appendix A Appendix
@@
 A.2 More Details of Results
-Table 1 summarizes the average accuracy and standard deviation across ten sampled subsets.
-Table 1: Mean accuracy (with standard deviation) across 10 subsamples for each reranking method.
-Method Mean Accuracy Std. Dev.
-Coder 0.48 0.059
-CoderReviewer 0.51 0.043
-CodeRSA 0.55 0.044
-random 0.43 0.030
-As shown in Table 1, CodeRSA achieves the highest average accuracy (0.55), outperforming Coder (0.48), CoderReviewer (0.51), and the random baseline (0.43).
-This indicates that CodeRSA provides stronger overall performance compared to the other methods.
-From a standard-deviation perspective, CoderReviewer and CodeRSA both exhibit relatively stable performance across different subsamples.
-By contrast, Coder has the highest standard deviation (0.059), suggesting greater sensitivity to varying combinations of coding problems.
-Figure 5: Performance difference between CodeRSA and CoderReviewer.
-Fig.
-5 depicts the per-trial accuracy difference between CodeRSA and CoderReviewer (i.e., CodeRSA accuracy minus CoderReviewer accuracy).
-Bars above zero indicate that CodeRSA outperforms CoderReviewer, while the single negative bar (Trial 2) represents the only instance where CodeRSA yields a lower accuracy.
-In most trials, CodeRSA demonstrates gains of up to 10%, highlighting its consistent advantage.
-A.3 Ablation Study
-Figure 6: Box plot of accuracy in ablation study.
-To investigate the impact of the log transformation in the calculation of the pragmatic speaker score, we conducted an ablation study comparing the CodeRSA’s modeling of pragmatic speaker score:
-$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})}\;,$
-$\text{R}_{S_{1}}(i\mid c)\;\in\;(-1,\,0).$
-against a variant that directly utilizes the normalized probability:
-$\text{P}_{S_{1}}(i\mid c)=\frac{\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})},$
-referred to as CodeRSA_nolog.
-Fig.
-6 displays the box plots of accuracy for different methods on ten subsets.
-From the results, CodeRSA achieves the highest mean accuracy with a narrow interquartile range, indicating stable performance.
-In contrast, CodeRSA_nolog shows a marked decrease in accuracy, evidenced by a lowest mean and a compressed interquartile range.
-This observation aligns with previous findings by Schuster et al.
-(2024), which reported that the basic RSA model underperforms the baseline on their SpreadNaLa dataset.
-These results suggest that applying a logarithmic transformation to expand the probability intervals can substantially enhance overall performance.
-A.4 A Geometric Mean Perspective on the Log-Normalization Formula
-In our proposed CodeRSA model, we define the pragmatic speaker score as follows:
-$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime})}\;,$
-$\text{R}_{S_{1}}(i\mid c)\;\in\;(-1,\,0).$
-where $\text{P}_{L_{0}}(c\mid i)$ is the probability estimated by the literal listener for candidate code $c$ given instruction $i$, and $I$ denotes the set of potential instructions.
-The geometric mean of the set $\{\text{P}_{L_{0}}(c\mid i^{\prime}):i^{\prime}\in I\}$ is defined as:
-$\text{GM}=\left(\prod_{i^{\prime}\in I}\text{P}_{L_{0}}(c\mid i^{\prime})\right)^{\frac{1}{|I|}},$
-and its logarithm is given by:
-$\log\text{GM}=\frac{1}{|I|}\sum_{i^{\prime}\in I}\log\text{P}_{L_{0}}(c\mid i^{\prime}).$
-Thus:
-$\text{R}_{S_{1}}(i\mid c)=-\frac{\log\text{P}_{L_{0}}(c\mid i)}{|I|\cdot\log\text{GM}}\;.$
-In this sense, normalizing the log probability for a particular instruction $i$ by the sum over all potential instructions can be seen as a way of capturing its relative contribution in a manner analogous to the geometric mean.
+In this subsection, we provide further results on different datasets and models to further validate the robustness of CodeRSA.
+Specifically, we evaluate on the HumanEval dataset as well as on MBPP, using both Llama-3-8B-Instruct and Qwen2.5-7B-Instruct.
+Figure 5: Accuracy of CodeRSA on HumanEval with Qwen2.5-7B-Instruct. The shaded region shows the stable band.
+Figure 6: Accuracy of CodeRSA on HumanEval with Llama-3-8B-Instruct. The shaded region shows the stable band.
+Figure 7: Accuracy of CodeRSA on MBPP with Qwen2.5-7B-Instruct. The shaded region shows the stable band.
+Across all settings, several consistent trends can be observed.
+Clustering effectiveness. Removing the clustering step leads to a noticeable drop in accuracy, highlighting its role in reducing redundancy and stabilizing pragmatic reasoning.
+Calibration robustness. CodeRSA is not strongly sensitive to the calibration parameter $\alpha$; performance remains stable across a relatively wide range rather than relying on a finely tuned value.
+Superior accuracy.
+CodeRSA consistently achieves higher accuracy than both baselines.
+On the HumanEval dataset, we observe some fluctuations in performance, and the overall accuracy is relatively low.
+This may be partly due to randomness or parameter settings in the experiments.
+However, since all reranking methods are evaluated under the same conditions, the relative comparison between them remains fair and informative.
+These findings confirm that the improvements achieved by CodeRSA are reliable across different models and datasets.
+The calibration parameter $\alpha$ is shown to be both interpretable and stable, further supporting the practicality of the approach.
+A.3 Accuracy vs. Number of Sampled Candidates (MBPP)
+Figure 8: MBPP: Mean accuracy vs.
+number of randomly sampled candidates per problem ($n$).
+Curves show CodeRSA (orange), Coder (green), and CoderReviewer (purple).
+Shaded regions indicate the standard deviation across multiple random samplings.
+To further examine the impact of candidate diversity on reranking performance, we conducted a controlled study varying the number of sampled code candidates per MBPP problem ($n=1\ldots 10$).
+For each value of $n$, we randomly sampled $n$ candidates from the pool of ten generated solutions and applied three reranking strategies: Coder, CoderReviewer, and our proposed CodeRSA.
+The experiment was repeated ten times with different random seeds for each value of $n$, and the figure reports the mean accuracy and its standard deviation across runs.
+When $n=1$, all methods yield identical results, as no reranking can occur.
+As $n$ increases, performance improves for all methods due to a broader candidate set, but the gain plateaus after approximately $n=7$.
+Across all sampling levels, CodeRSA achieves the highest accuracy, maintaining a margin of roughly 2–3 percentage points over CoderReviewer and up to 5 points over Coder.
+This shows that pragmatic reasoning allows CodeRSA to better leverage candidate diversity while remaining robust to sampling variability.
+The narrow confidence bands further indicate stable performance even under random candidate selection, confirming its reliability when generation stochasticity varies across runs or models.
+In our main experiments, we fixed $n=10$ candidates per problem as a practical balance between computational cost and runtime.
+The results here further suggest that model performance is not strongly dependent on candidate set size.
+Future work could explore larger candidate pools when computational resources permit.
+A.4 An example of clustering
+In the following presentation, each item is denoted in the format:
+$\textbf{code\_X}:\text{Instruction generated from this code}$
+This means that the left-hand side (code_X) represents the identifier of the function implementation, and the right-hand side is the instruction generated based on it.
+Main Cluster: Maximum value with frequency condition
+  • code_1: Create a function that takes a list of numbers. Returns the largest number that appears at least as many times as its value in the list. If no such number exists, return -1.
+  • code_6: Create a function that takes a list of integers and returns the maximum value that appears at least as many times as its value. If no such value exists, return -1.
+  • code_8: Create a function that takes a list of numbers and returns the maximum integer that occurs at least as many times as its value.
+If multiple such numbers exist, return the largest one.
+If no such number exists, return -1.
+Cluster 2: Most frequent element
+  • code_5: Create a function that takes a list of integers.
+Returns the number that appears most frequently in the list.
+If there are multiple such numbers with the same frequency, return the largest one.
+  • code_10: Create a function that takes a list of integers.
+Returns the most frequent integer greater than 0.
+If multiple integers have the same highest frequency, return the smallest one.
+If the list is empty, return -1.
+Cluster 3: Repeated integers
+  • code_4: Create a function that takes a list of integers and returns the smallest positive integer that appears more than once. If no such integer exists, return -1.
+  • code_7: Create a function that takes a list of integers and returns the maximum value that appears more than once. If no such value exists, return -1.
+Cluster 4: Missing positive integer
+  • code_3: Create a function that takes a list of integers and returns the first missing positive integer. If the list is empty, return -1.
+Cluster 5: First/last integer with frequency condition
+  • code_2: Create a function that takes a list of integers and finds the first integer that occurs at least as many times as its value. If no such integer is found, return None.
+  • code_9: Create a function that takes a list of integers. Returns the first integer that appears at least as many times as its value in the list. If no such integer exists, return None.
 A.5 Prompt Used

### v4

@@
 Abstract
-Large Language Models (LLMs) have demonstrated impressive potential in translating natural language (NL) instructions into program code.
-However, user instructions often contain inherent ambiguities, making it challenging for LLMs to generate code that accurately reflects the user’s true intent.
-To address this challenge, researchers have proposed approaches that produce multiple candidates of the program code and then rerank them to identify the best solution.
-In this paper, we propose CodeRSA, a novel code candidate reranking mechanism built upon the Rational Speech Act (RSA) framework, designed to guide LLMs toward more comprehensive pragmatic reasoning about user intent.
-We evaluate CodeRSA using Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct on two widely used code generation benchmarks, HumanEval and MBPP.
-Our experiment results show that CodeRSA consistently outperforms common baselines, surpasses the state-of-the-art approach in most cases, and demonstrates robust overall performance.
-These findings underscore the effectiveness of integrating pragmatic reasoning into code candidate reranking, offering a promising direction for enhancing code generation quality in LLMs.
+Pragmatic reasoning is pervasive in human–human communication — it allows us to leverage shared knowledge and counterfactual reasoning in order to infer the intention of a conversational partner given their ambiguous or underspecified message.
+In human–computer communication, underspecified messages often represent a major challenge: for instance, translating natural language instructions into code is difficult when user instructions contain inherent ambiguities.
+In the present paper, we aim to scale up the pragmatic “Rational Speech Act” framework to naturalistic language-to-code problems, and propose a way of dealing with multiple meaning-equivalent instruction alternatives, an issue that does not arise in previous toy-scale problems.
+We evaluate our method, CodeRSA, with two recent LLMs (Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct) on two widely used code generation benchmarks (HumanEval and MBPP).
+Our experimental results show that CodeRSA consistently outperforms common baselines, surpasses the state-of-the-art approach in most cases, and demonstrates robust overall performance.
+Qualitative analyses demonstrate that it exhibits the desired behaviour for the right reasons.
+These findings underscore the effectiveness of integrating pragmatic reasoning into a naturalistic complex communication task, language-to-code generation, offering a promising direction for enhancing code generation quality in LLMs and emphasizing the importance of pragmatic reasoning in complex communication settings.
 Pragmatic Reasoning improves LLM Code Generation
@@
 However, given the intrinsic complexities of coding and the potential ambiguities in user input, producing code in a single attempt may fail to explore the vast solution space, overlooking correct or higher-quality solutions (Liu et al., 2024a).
-A standard practice to address this shortcoming is to sample multiple solutions, which we refer to as code candidates (Chen et al., 2021; Brown et al., 2024), and to rerank them.
-Researchers have proposed various reranking strategies for code candidates, broadly divided into execution-driven and content-driven approaches.
-Due to the unreliability of automatically generated test suites (Chen et al., 2022) and the potential safety risks associated with code execution (Yetiştiren et al., 2023), we focus instead on content-driven methods, which evaluate the generated text, often relying on token-level probabilities.
-For example, Coder Reranking scores each candidate based on the cumulative probability of its tokens, sometimes however favoring ‘degenerate solutions” (generic or repetitive code) with disproportionately high token probabilities (Zhang et al., 2023a).
+A standard practice to address this shortcoming is to sample multiple solutions, which we refer to as code candidates (Chen and others, 2021; Brown et al., 2024), and to rerank them.
 Figure 1: A comparison of our approach CodeRSA (top) compared to CoderReviewer (bottom).
-When viewing code generation as a communicative process in which an LLM listens to the user’s intentions (Ouyang et al., 2022), Coder Reranking evaluates candidate solutions solely from the listener’s perspective.
-Yet, research on human communication suggests that effective listeners reason about the speaker (who in turn reasons about the listener) Grice (1975).
-Frank and Goodman (2012; 2016) provided a principled method for quantifying this process based on a probabilistic framework rooted in game-theoretic notions, called the Rational Speech Act (RSA) framework.
-Pu et al.
+When viewing code generation as a communicative process in which an LLM listens to the user’s intentions (Ouyang et al., 2022), the different code candidates represent the different possible user intents and the instruction represents an ambiguous or underspecified message.
+Pragmatic reasoning then involves considering not only the different possible intended meanings, but also counterfactual reasoning about what instruction the speaker would have given if they had a specific intent.
+These ideas go back to Grice (1975) and have been mathematically formalized using a probabilistic model rooted in game-theoretic notions, as the "Rational Speech Act" (RSA) framework in Frank and Goodman (2012; 2016).
+Code reranking can thus be performed with respect to the RSA reasoning process.
+In previous research, Pu et al.
 (2020, 2024) demonstrated the effectiveness of the RSA framework for program generation in a simple domain (regular-expression synthesis), while Schuster et al.
@@
 1, Pragmatic Speaker Part, where the main cluster refers to the one containing the original description $I_{0}$).
-We conducted experiments using CodeRSA with Llama-3-8B-Instruct, one of the latest language models from the Llama family (Grattafiori et al., 2024), and Qwen-2.5-7B-Instruct, a recent instruction-tuned model from the Qwen series (Yang et al., 2024), on two widely used code generation benchmarks: OpenAI’s HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021).
-Our experimental results demonstrate that CodeRSA reliably outperforms the Coder and CoderReviewer Reranking methods.
+We conducted experiments using CodeRSA with Llama-3-8B-Instruct, one of the latest language models from the Llama family (Grattafiori and others, 2024), and Qwen-2.5-7B-Instruct, a recent instruction-tuned model from the Qwen series (Yang et al., 2024), on two widely used code generation benchmarks: OpenAI’s HumanEval (Chen and others, 2021) and MBPP (Austin et al., 2021).
+Our experimental results demonstrate that CodeRSA reliably outperforms the simpler Coder and CoderReviewer Reranking methods.
 Our qualitative analysis reveals how the CodeRSA enables better candidate selection, promoting a more comprehensive understanding of user intent.
@@
 Previous research has extensively explored generating code from natural language using neural network models (Ling et al., 2016; Rabinovich et al., 2017; Hayati et al., 2018).
-Recently, large language models (LLMs) have propelled significant advances in this area, driven by the transformer Vaswani (2017) architecture and large-scale pretraining.
+Recently, large language models (LLMs) have propelled significant advances in this area, driven by the transformer Vaswani et al.
+(2017) architecture and large-scale pretraining.
 Their performance on code generation tasks often surpasses that of traditional models, and in many cases even rivals human programmers (Ni et al., 2024; Becker et al., 2023).
@@
 Code Reranking Methods.
+Researchers have proposed various reranking strategies for code candidates, broadly divided into execution-driven and content-driven approaches.
 Execution-driven reranking methods such as CodeT (Chen et al., 2022) and AgentCoder (Huang et al., 2024) evaluate code candidates by running them against automatically generated test suites.
@@
 In contrast, content-driven reranking methods are far more versatile because they do not rely on execution and are not even confined to coding tasks.
+One such model is Coder Reranking (Chen and others, 2021).
 Coder Reranking.
-Chen et al.
-(2021) rerank code candidates by estimating $\text{P}(c\mid i)$, where $c$ denotes the generated code candidate and $i$ denotes the given instruction.
+Chen and others (2021) rerank code candidates by estimating $\text{P}(c\mid i)$, where $c$ denotes the generated code candidate and $i$ denotes the given instruction.
 This process can also be called Coder Reranking because the LLM is a mere Coder that estimates the candidate probability based on the corresponding instruction.
@@
 In our experiments, we treat $\alpha$ as a tunable hyperparameter and find that performance is stable across a broad range of values (see Section 5).
-Clustering Paraphrases.
-While the basic RSA formulation operates directly over the instruction set $I$, it can suffer from over-interpreting superficial variations in wording when many instructions are semantically equivalent and differ only in surface form.
-In such cases, RSA allocates probability mass across paraphrases as if they were meaningful distinct alternatives, diluting the signal and reducing accuracy.
+Clustering Paraphrases. While the basic RSA formulation operates directly over the instruction set $I$,
 Semantically equivalent instructions:
@@
 Table 1: Examples of equivalent and non-equivalent instructions from MBPP. The first group expresses the same semantics, while the second differs in meaning.
-To mitigate this, CodeRSA employs a semantic clustering.
+it can suffer from over-interpreting superficial variations in wording when many instructions are semantically equivalent and differ only in surface form.
+In such cases, RSA allocates probability mass across paraphrases as if they were meaningful distinct alternatives, diluting the signal and reducing accuracy.
+To mitigate this, CodeRSA employs semantic clustering.
 Candidate instructions are grouped into semantic clusters $\mathcal{C}=\{C_{1},\ldots,C_{K}\}$ using an LLM-based equivalence test (implementation details in Section 4), so that pragmatic reasoning operates over clusters rather than individual instructions.
@@
 We evaluate on two widely used code generation benchmarks.
-HumanEval (Chen et al., 2021) contains 164 Python programming problems, each presented as an unfinished function with a natural language instruction.
-MBPP (Austin et al., 2021) includes 257 short programming tasks with natural language prompts.
+HumanEval (Chen and others, 2021) contains 164 Python programming problems, each presented as an unfinished function with a natural language instruction.
+MBPP (Austin et al., 2021) includes 974 short programming tasks with natural language prompts.
 HumanEval offers balanced difficulty, while MBPP introduces greater lexical variety.
@@
 A sensitivity check with varying numbers of sampled candidates is provided in Appendix A.4.
-For our experiments, we use Llama-3-8B-Instruct (Grattafiori et al., 2024) and Qwen-2.5-7B-Instruct (Yang et al., 2024), two instruction-tuned LLMs of comparable scale.
+For our experiments, we use Llama-3-8B-Instruct (Grattafiori and others, 2024) and Qwen-2.5-7B-Instruct (Yang et al., 2024), two instruction-tuned LLMs of comparable scale.
 Llama-3-8B-Instruct balances efficiency with strong generation quality, while Qwen-2.5-7B-Instruct provides a competitive open-source alternative.
@@
 Future work could explore hybrid setups, for example, using coder models for program synthesis combined with general-purpose instruction-tuned models for reasoning about instructions.
+Our goal is to examine whether RSA-style pragmatic reasoning can be effectively applied to realistic code generation through appropriate methodological design.
+We evaluate our approach on two established benchmarks (MBPP and HumanEval) using mid-sized instruction-tuned models, which provide sufficient headroom for reranking.
+For example, Llama-3-8B-Instruct improves from 51.3% pass@1 to 78.2% pass@10 on MBPP, indicating that correct solutions are often generated but not reliably selected in a single attempt.
+In such settings, the effect of reranking is most clearly observable.
+Appendix A.3 further reports a sanity check with Llama-3-70B-Instruct Grattafiori and others (2024), confirming that CodeRSA behaves consistently when applied to a substantially larger model.
 4.2 Implementation of Reranking Methods
@@
 5.1 Quantitative Analysis
+Figure 3: Accuracy of CodeRSA across different values of the calibration parameter $\alpha$. The shaded region indicates a stable performance band.
+(a) Details of question and two generated examples
+(b) Coder Score Comparison
+(c) Pragmatic Score Comparison
+Figure 4: Qualitative Example: Bias in Coder vs. CodeRSA Correction
 In this section, we analyze the quantitative performance of CodeRSA with respect to the calibration parameter $\alpha$ (defined in Section 3), which controls the influence of the prior through temperature scaling, using the MBPP dataset and Llama-3-8B-Instruct model.
-Figure 3: Accuracy of CodeRSA across different values of the calibration parameter $\alpha$. The shaded region indicates a stable performance band.
 Here, accuracy denotes the proportion of test instances where the reranking method selects a candidate that passes all test cases provided by the benchmark.
@@
 These results demonstrate that CodeRSA’s pragmatic reasoning, enhanced by clustering, is not overly sensitive to calibration.
-We report further results covering different models (Qwen2.5-7B-Instruct) and datasets (HumanEval) in Appendix A.2, which confirm the same overall trends.
-(a) Details of question and two generated examples
-(b) Coder Score Comparison
-(c) Pragmatic Score Comparison
-Figure 4: Qualitative Example: Bias in Coder vs. CodeRSA Correction
+We report further results covering different models (Qwen2.5-7B-Instruct) and datasets (HumanEval) in Appendix A.2, which confirm the same pattern of CodeRSA consistently improving performance over baseline methods.
 5.2 Qualitative Analysis
@@
 Future work should investigate variable cost structures to better capture these nuances.
-In Section 4, we argued that CodeRSA, as a reranking approach, is most beneficial in situations where the dataset is not too easy (when a simple Coder model already achieves ceiling performance) and not too difficult, such that we can still obtain a high quality probability distribution over instructions and over code candidates.
-This raises the question of the relevance of pragmatic reasoning for code generation, and more generally in communication.
+In Section 4, we argued that CodeRSA is most beneficial in situations where the dataset is not too easy (when a simple Coder model already achieves ceiling performance) and not too difficult, i.e., good candidates need to be contained in the set of sampled candidates which is then re-ranked.
+That is, an important pre-condition for the applicability of the CodeRSA approach is the ability to obtain a high quality probability distribution over instructions and over code candidates.
+We argue that in complex tasks like language instructions for generating code, a situation where the correct candidate is in the set of sampled candidates, but where it is not ranked first, is not just due to bad model performance that will become better with improved models, but is expected to be a systematic phenomenon that is due to the inherent ambiguity and underspecification in human natural language instructions.
+Dealing with this type of ambiguity is not a matter of better language modelling, but inherently requires the counterfactual reasoning process implemented in RSA models.
 Research on human communication has demonstrated the importance of pragmatic reasoning, even though it introduces additional computational overhead.
-At the same time, studies suggest that humans may rely on simple heuristics or amortized estimates (Pu et al., 2024), avoiding iterative reasoning in easy cases while still engaging in full pragmatic reasoning when tasks are more complex.
+At the same time, studies suggest that humans may switch between different strategies where they sometimes rely on the expensive reasoning process, and sometimes rely on simple heuristics or amortized estimates (Lieder and Griffiths, 2020; Tscshantz et al., 2023; Dasgupta et al., 2018; Pu et al., 2024), avoiding iterative reasoning in easy cases while still engaging in full pragmatic reasoning when tasks are more complex.
+Such strategies can also be explored in future CodeRSA research.
 7 Conclusion
-This work introduces CodeRSA, a candidate reranking algorithm for the generation of program code grounded in the Rational Speech Act framework.
+This work introduces CodeRSA, a pragmatics-inspired approach to dealing with ambiguity and underspecification in human instructions for code generation.
 By modeling the iterative reasoning of a pragmatic listener about a pragmatic speaker, CodeRSA consistently outperforms the Coder Reranking baseline and surpasses the state-of-the-art CoderReviewer approach.
-A qualitative analysis further reveals that, even when incorporating certain idealized assumptions and variations, CodeRSA remains faithful to the core principles of the RSA framework.
+Our experiments showed highly consistent results for two models from different model families and two different benchmarks.
+A qualitative analysis further reveals that, even when incorporating certain idealized assumptions and variations, CodeRSA remains faithful to the core principles and goals of the RSA framework.
 These results highlight the effectiveness of applying well-established linguistic frameworks to enhance reasoning in language models, opening new avenues for research and development in code-related tasks.
+Some of our insights related to practical and naturalistic application domains, such as the need for clustering of very similar instructions or the need to re-weight probability estimates for instructions vs.
+meaning representations may also be applicable to human-computer interaction tasks beyond language-to-code.
 8 Limitations
@@
 Such improvements would allow CodeRSA to scale more effectively and broaden its applicability to larger code generation tasks.
-Another limitation is that, although our experiments already cover two models (Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct) and two datasets (HumanEval and MBPP), the scope remains relatively narrow.
-We are currently working on incorporating additional balanced-difficulty datasets such as DS-1000 (Lai et al., 2023)), along with further open-source models like Mistral (Jiang et al., 2023) and newer Qwen releases beyond Qwen-2.5.
+Another limitation concerns the scope of models and benchmarks considered.
+In the current study, we deliberately focus on mid-sized models and well-established benchmarks as a proof-of-concept design choice, enabling fair and clearly observable comparisons with existing reranking baselines, rather than aiming for exhaustive coverage across model scales or benchmark families.
+As a result, although our experiments already cover two models (Llama-3-8B-Instruct and Qwen-2.5-7B-Instruct) and two datasets (HumanEval and MBPP), the overall scope remains relatively narrow.
+In future work, we plan to incorporate additional balanced-difficulty datasets such as DS-1000 (Lai et al., 2023), along with further open-source models like Mistral (Jiang et al., 2023) and newer Qwen releases beyond Qwen-2.5.
 This expansion will allow us to evaluate reranking methods across a wider range of scenarios, ultimately leading to a more comprehensive assessment of our approach.
+At the same time, ensuring an appropriate match between model capability and benchmark difficulty remains an important challenge for designing informative evaluations.
 Appendix A Appendix
-A.1 A Conjecture: Explaining CoderReviewer from an RSA Perspective
-In the RSA framework, a pragmatic listener’s posterior over a candidate $c$ given an instruction $i$ is commonly expressed as:
-$\text{P}_{L_{1}}(c\mid i)\propto\text{P}_{S_{1}}(i\mid c)\cdot\text{P}(c),$
-where $\text{P}_{S_{1}}(i\mid c)$ represents how likely a pragmatic speaker would be to produce instruction $i$ when the correct candidate is $c$, and $\text{P}(c)$ is the prior likelihood of $c$.
-Translating this perspective to LLMs, we hypothesize that when generating instructions (the “Reviewer” role), it is relatively straightforward for the model to produce abstract instructions from concrete code.
-Since code is unambiguous, the LLM can approximate a pragmatic speaker:
-$\text{P}_{\text{LLM}}(i\mid c)\approx\text{P}_{S_{1}}(i\mid c).$
-However, generating code from abstract instructions (the “Coder” role) is substantially more difficult.
-In this setting, the LLM may effectively revert to estimating a prior over possible candidates, thereby approximating:
-$\text{P}_{\text{LLM}}(c\mid i)\approx\text{P}(c).$
-From this RSA standpoint, the CoderReviewer paradigm can be considered a simplified, yet broad, modeling of a pragmatic listener.
+A.1 A Conjecture: Viewing CoderReviewer through an RSA Lens
+CoderReviewer reranks candidates using the bidirectional criterion
+$\text{P}_{\text{CR}}(c\mid i_{0})\ \propto\ \text{P}_{\text{LLM}}(c\mid i_{0})\cdot\text{P}_{\text{LLM}}(i_{0}\mid c),$
+which rewards candidates that are (i) likely under the instruction (Coder) and (ii) able to “explain” the instruction when the code is given as context (Reviewer).
+In RSA, a pragmatic listener instead ranks candidates according to
+$\text{P}_{L_{1}}(c\mid i_{0})\ \propto\ \text{P}_{S_{1}}(i_{0}\mid c)\cdot\text{P}(c),$
+combining (a) a speaker term that captures how well the instruction identifies the intended candidate among alternatives, and (b) a prior over candidates.
+We conjecture that CoderReviewer can be interpreted as a lightweight, heuristic approximation to this RSA listener objective on the restricted candidate set used for reranking.
+Intuitively, producing a faithful natural-language description from a concrete piece of code is comparatively constrained, so the reverse model $\text{P}_{\text{LLM}}(i_{0}\mid c)$ may behave like an amortized proxy for a speaker preference $\text{P}_{S_{1}}(i_{0}\mid c)$ (without explicitly normalizing over an alternative-instruction set or modeling instruction costs).
+Meanwhile, when $i_{0}$ is underspecified, the forward term $\text{P}_{\text{LLM}}(c\mid i_{0})$ can be strongly influenced by instruction-independent code preferences (well-formedness, idiomatic patterns, and length/format biases), and thus acts as a task-local proxy for the candidate prior $\text{P}(c)$ on the finite set of sampled candidates.
+Under this view, multiplying the two terms resembles combining a speaker-like compatibility signal with a prior-like plausibility signal, as in RSA.
 A.2 More Details of Results
@@
 The calibration parameter $\alpha$ is shown to be both interpretable and stable, further supporting the practicality of the approach.
-A.3 Accuracy vs. Number of Sampled Candidates (MBPP)
+A.3 Additional Experiment: Larger-Model Sanity Check on MBPP
+As an additional sanity check, we evaluate CodeRSA with a larger base model.
+Since RSA-style reranking is computationally expensive, we keep this experiment limited in scope.
+We use Llama-3-70B-Instruct Grattafiori and others (2024) on MBPP and restrict evaluation to instances that are unsolved at pass@1 but solved at pass@10, resulting in 84 problems where reranking is non-trivial.
+Method Solved Acc.
+Coder 33/84 39.3%
+CoderReviewer 41/84 48.8%
+CodeRSA (best $\alpha=0.9$) 43/84 51.2%
+CodeRSA (avg. $\alpha\in[0.5,1.5]$) 35.5/84 42.3%
+Table 2: Sanity check on Llama-3-70B-Instruct for MBPP, evaluated on 84 instances that are unsolved at pass@1 but solved at pass@10.
+For CodeRSA, “best” is the best value in a sweep with step size 0.1; “avg.” is the mean over $\alpha\in\{0.5,0.6,\ldots,1.5\}$.
+Table 2 reports the number of solved problems (out of 84) and the corresponding accuracy after reranking.
+CodeRSA achieves the best result at $\alpha=0.9$ in our sweep (step size 0.1), outperforming Coder and slightly improving over CoderReviewer on this subset.
+Averaged over $\alpha\in[0.5,1.5]$, CodeRSA remains above Coder, indicating that scaling to a larger model does not lead to abnormal behavior, while calibration still matters for peak performance.
+A.4 Accuracy vs. Number of Sampled Candidates (MBPP)
+To further examine the impact of candidate diversity on reranking performance, we conducted a controlled study varying the number of sampled code candidates per MBPP problem ($n=1\ldots 10$).
+For each value of $n$, we randomly sampled $n$ candidates from the pool of ten generated solutions and applied three reranking strategies: Coder, CoderReviewer, and our proposed CodeRSA.
+The experiment was repeated ten times with different random seeds for each value of $n$, and the figure reports the mean accuracy and its standard deviation across runs.
 Figure 8: MBPP: Mean accuracy vs.
@@
 Shaded regions indicate the standard deviation across multiple random samplings.
-To further examine the impact of candidate diversity on reranking performance, we conducted a controlled study varying the number of sampled code candidates per MBPP problem ($n=1\ldots 10$).
-For each value of $n$, we randomly sampled $n$ candidates from the pool of ten generated solutions and applied three reranking strategies: Coder, CoderReviewer, and our proposed CodeRSA.
-The experiment was repeated ten times with different random seeds for each value of $n$, and the figure reports the mean accuracy and its standard deviation across runs.
 When $n=1$, all methods yield identical results, as no reranking can occur.
@@
 Future work could explore larger candidate pools when computational resources permit.
-A.4 An example of clustering
+A.5 An example of clustering
 In the following presentation, each item is denoted in the format:
@@
   • code_9: Create a function that takes a list of integers. Returns the first integer that appears at least as many times as its value in the list. If no such integer exists, return None.
-A.5 Prompt Used
-A.5.1 For Generating the Additional Instruction:
+A.6 Prompt Used
+A.6.1 For Generating the Additional Instruction:
 ##Write an instruction for given python function##
@@
 ###instruction start###
-A.5.2 For Calculating the Reviewer Score (An Example):
+A.6.2 For Calculating the Reviewer Score (An Example):
 def any_int(x, y, z):